### v1

# Which Reasoning Trajectories Teach Students to Reason Better? A Simple Metric of Informative Alignment

Abstract
Long chain-of-thought (CoT) trajectories provide rich supervision signals for distilling reasoning from teacher to student LLMs. However, both prior work and our experiments show that trajectories from stronger teachers do not necessarily yield better students, highlighting the importance of data–student suitability in distillation. Existing methods assess suitability primarily through student likelihood, favoring trajectories that closely align with the model’s current behavior but overlooking more informative ones. Addressing this, we propose Rank-Surprisal Ratio (RSR), a simple metric that captures both alignment and informativeness to assess the suitability of a reasoning trajectory. RSR is motivated by the observation that effective trajectories typically combine low absolute probability with relatively high-ranked tokens under the student model, balancing learning signal strength and behavioral alignment. Concretely, RSR is defined as the ratio of a trajectory’s average token-wise rank to its average negative log-likelihood, and is straightforward to compute and interpret. Across five student models and reasoning trajectories from 11 diverse teachers, RSR strongly correlates with post-training performance (average Spearman 0.86), outperforming existing metrics. We further demonstrate its practical utility in both trajectory selection and teacher selection.
$\dagger$ Corresponding authors. Inquiries may be sent to: yumingyang23@m.fudan.edu.cn, zouyicheng@pjlab.org.cn, qz@fudan.edu.cn* Code is available at https://github.com/UmeanNever/RankSurprisalRatio.
1 Introduction
Recent advances in reasoning-oriented large language models (LLMs) are largely driven by their ability to generate long chain-of-thought (CoT) trajectories [12, 48]. Beyond enabling complex inference at test time, such trajectories also provide powerful supervision signals for training student models [35, 28] or cold-starting reinforcement learning [39] through supervised fine-tuning (SFT).
Yet, stronger reasoning teachers do not necessarily yield better students [22, 11]. Our extensive experiments show that the post-training effectiveness of reasoning trajectories varies substantially across student models, indicating that the suitability between data and student is critical for effective learning. Existing data engineering methods assess data suitability primarily through the student’s probability assignments [47, 18], favoring high-likelihood trajectories that align closely with the model’s current behavior. Such trajectories, however, often provide limited new learning signals. In contrast, more informative trajectories are typically less familiar to the student and thus overlooked by these methods. This leads to a fundamental Informative Alignment challenge: how to identify reasoning data that are both well aligned with the student and sufficiently informative?
To address this challenge, we propose a simple yet effective metric, Rank-Surprisal Ratio (RSR), which quantifies the suitability of a reasoning trajectory for a given student by jointly capturing alignment and informativeness. Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting absolute unfamiliarity and relative familiarity. Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while still containing behavioral patterns within the student’s prior experience, such that their tokens rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
Figure 1: Illustration of the intuition behind RSR. Suitable reasoning trajectories should balance informativeness and alignment by having low absolute probability but relatively high-ranked tokens under the student model.
Based on this insight and consistent numerical patterns observed in simulation studies, we define our suitability metric, Rank-Surprisal Ratio, as the ratio between a trajectory’s average token-wise rank [^1] Higher-ranked tokens have lower rank values. and its average negative log-likelihood (surprisal). RSR can be computed with a single forward pass, requires no additional verifier or test data, and is straightforward to interpret. Lower RSR indicates better informative alignment, identifying trajectories that are both informative and well aligned with the student.
We validate the effectiveness of Rank-Surprisal Ratio through correlation analyses on 5 student LLMs using math reasoning trajectories generated by 11 representative teacher models. Across all students, the RSR of trajectories exhibits a strong correlation with post-training performance, achieving an average Spearman correlation of 0.86 and consistently outperforming alternative metrics. Furthermore, to explore its practical value in data engineering, we apply RSR to trajectory selection and teacher selection. Our experiments show that RSR not only selects more effective training trajectories for each problem from candidates generated by diverse teachers, but also identifies more suitable teacher models using only a small amount of data, consistently outperforming existing selection methods across all five students in both settings.
Our main contributions are three-fold:
  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data–student suitability.
  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
  • We demonstrate the practical utility of Rank-Surprisal Ratio in two data engineering scenarios, trajectory selection and teacher selection, where it serves as an effective criterion and outperforms existing methods.
2 The Need for Student-Specific Data
To understand which types of reasoning trajectories most effectively improve student models after SFT, we conduct a comprehensive large-scale study involving five widely adopted student models and eleven diverse reasoning-oriented teacher models, yielding 55 teacher–student pairings. We perform SFT experiments for each of these pairs.
Teacher Models Params Student Models (Base) Teacher Performance
Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B
Deepseek-R1 671B 77.1 28.1 47.3 55.8 29.6 91.1
Qwen-3-235B-Thinking 235B 71.8 22.0 45.0 53.4 26.4 91.2
GPT-OSS-120B 120B 66.7 15.2 40.7 47.9 22.9 88.3
Nemotron-Super 49B 72.2 23.7 48.3 56.4 33.0 82.3
QwQ-32B 32B 77.4 27.1 52.0 61.2 33.0 85.2
Qwen-3-30B-Thinking 30B 77.2 26.7 50.0 58.8 31.2 92.3
Magistral-Small 24B 68.8 22.8 47.6 52.2 30.6 71.0
GPT-OSS-20B 20B 69.5 17.9 42.7 48.4 24.4 83.4
Phi-4-Reasoning-Plus 14B 54.1 14.5 35.2 40.2 18.2 72.7
Qwen-3-8B 8B 74.6 26.5 52.0 61.2 34.2 82.5
Qwen-3-4B-Thinking 4B 76.8 28.2 51.8 61.9 33.3 87.3
Table 1: Distillation results showing post-training reasoning performance of student models trained on trajectories from different teacher models, evaluated by average Acc@4 on AIME’25, AIME’24, AMC’23, and MATH500. Darker and lighter shading indicate the best and second-best results, respectively. Student performance varies significantly across teacher–student pairs, highlighting the importance of data–student suitability.
2.1 Experimental Settings
Our teacher-student pairing study involves two major steps: (1) For each teacher model, we prompt it to generate a long CoT response for each math problem in our 5000-problem set (see § A.1), forming a trajectory dataset specific to that teacher. (2) For each teacher–student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance. All students are pretrained base models. To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher–student pair. Reported results are averaged over these three runs. More implementation details are provided in the appendix.
Teachers
We use 11 reasoning LLMs spanning 4B to 671B parameters across multiple model families, including GPT-OSS [3], DeepSeek [12], Qwen [39], LLaMA-Nemotron [4], and Phi [1].
Benchmarks
We evaluate the reasoning performance of the fine-tuned student models on four standard math benchmarks: AIME’25, AIME’24, AMC’23, and MATH500 [14] using the Acc@4 metric, and report results averaged across all benchmarks.
2.2 Results
Table 1 presents the results of our teacher–student pairing distillation study, revealing that:
Stronger teachers do not necessarily produce better students.
Teacher capability, whether measured by parameter scale or reasoning performance, does not reliably predict student improvement. For example, the 671B and 235B models often underperform smaller teachers such as QwQ-32B on multiple students. Similarly, teachers with strong reasoning performance do not consistently yield the best outcomes for all student models.
Data–student suitability is critical for eliciting reasoning improvements.
The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model. Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) may likewise be ineffective at improving stronger students (Qwen-3-14B). Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb. Overall, we find no simple teacher–student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
3 Measuring Data-Student Suitability
In this section, we explore metrics for measuring data–student suitability, with the goal of jointly capturing informativeness and alignment. We begin by introducing two fundamental token-level measures: surprisal and rank (§ 3.1), and analyzing the limitations of existing probability-based metrics (§ 3.2). We then abstract our insights on suitable reasoning data and conduct simulation studies to identify quantitative patterns (§ 3.3). Finally, we propose our trajectory-level metric (§ 3.4).
Teacher Models Student Performance $\uparrow$ Probability-based Metrics $\downarrow$ Rank-Surprisal Metrics $\downarrow$
[…]
Table 2: Comparison of the student’s post-training performance and data–student suitability metrics across trajectories from different teacher models, evaluated on Qwen-2.5-7B. Darker shading indicates higher performance or better suitability. Metrics whose trends align with performance (e.g., RSR) provide more reliable suitability estimates. Complete metric scores are provided in § C.2.
3.1 Surprisal and Rank
We introduce two different methods to quantify the amount of information carried by a token with respect to a student model. They serve as building blocks of our metric.
Surprisal (Negative Log-Likelihood)
A common measure is based on the probability of generating the current token $t_{k}$ given its preceding context $\mathbf{c}_{k}=(t_{1},\ldots,t_{k-1})$ under the student model $\theta$. For numerical stability, probabilities are typically transformed into the log space, and the negative log-likelihood is used as a measure of informativeness, also known as Surprisal [13]
$\mathrm{Surprisal}(t_{k})=-\log p_{\theta}\bigl(t_{k}\mid\mathbf{c}_{k}\bigr)$ (1)
Rank
Another method considers the rank of the current token within the model’s prediction distribution over the vocabulary $\mathcal{V}$. Formally, given the conditional distribution $p_{\theta}(\cdot\mid\mathbf{c}_{k})$, the rank of token $t_{k}$ is defined as the number of tokens with strictly higher probability [32].
$\mathrm{Rank}(t_{k})=1+\sum_{t^{\prime}\in\mathcal{V}}\mathbb{I}\!\left[p_{\theta}(t^{\prime}\mid\mathbf{c}_{k})>p_{\theta}(t_{k}\mid\mathbf{c}_{k})\right]$ (2)
Unlike surprisal, rank captures relative familiarity of the token by comparing target token against alternative candidates, revealing signals overlooked by probability-based measures. For instance, a token may receive low absolute probability while still being among the top-ranked candidates.
3.2 Limitations of Probability-Based Metrics
Existing work primarily relies on surprisal or log-probability to assess data suitability. For example, Zhang et al. [47] selects trajectories based on the average log-probability of response tokens under the student model. Since surprisal is the negation of log-probability, we implement this metric as the average surprisal, denoted as Avg-Surprisal. More recently, Just et al. [18] computes token-level log-probability based on a local context $\mathbf{c}_{k}^{\text{local}}$ (several preceding sentences), which we implement as average local surprisal (Avg-Surplocal). Under these metrics, trajectories with lower surprisal are considered more suitable.
However, as shown in Table 2, lower surprisal (i.e., higher probability) does not necessarily lead to better post-training reasoning performance. Both Avg-Surprisal and Avg-Surplocal assign lower surprisal to trajectories from Nemotron-Super and Magistral-Small, yet training on such trajectories fails to improve reasoning performance. Similar patterns are observed across all student models, indicating that probability-based metrics tend to favor data that are familiar but insufficiently informative.
At the other extreme, trajectories with very high surprisal (e.g., GPT-OSS-20B) also perform poorly. In contrast, trajectories with moderate surprisal values (e.g., Qwen-3-8B) achieve better results. This observation motivates us to investigate the mechanism underlying the surprisal trade-off.
3.3 Insight and Simulation
The above analysis suggests that effective teacher trajectories should balance the informativeness of the data with their alignment to the student’s current behavior: they should be neither overly similar to the student’s own generations nor excessively deviant from its predicted distribution.
At first glance, this balance may appear to pose a dilemma. However, we argue that it can be resolved by viewing informativeness and alignment through the lens of absolute unfamiliarity and relative familiarity. Informativeness does not require trajectories to be entirely unfamiliar; rather, it suffices that they deviate from the dominant patterns and thus have low absolute probability of being generated by the student. Conversely, alignment does not require exact agreement with the student’s outputs, but instead that the corresponding tokens have relatively higher likelihoods than other candidates in the vocabulary.
Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while still exhibiting behavioral patterns within the student’s prior experience. Therefore, tokens in such trajectories are assigned low absolute probability (high surprisal) by the student model while still ranking relatively high (low rank values) in its prediction distribution.
To validate these quantitative patterns and identify features that characterize effective learning trajectories, we conduct a simulation study.
Simulation Setting
We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we consider effective. Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$. The first mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant patterns, which arise from student’s massive general training data. The second mode $Z_{B}$, represents tokens whose patterns resemble strong reasoning trajectories, reflecting the student’s prior exposure to relevant data. We instantiate both $Z_{A}$ and $Z_{B}$ as Zipf distributions [52, 26] over $\mathcal{V}$. Then the student’s overall token-level prediction distribution $Z$ is constructed as a mixture of $Z_{A}$ and $Z_{B}$:
$Z=\pi\,Z_{A}+(1-\pi)\,Z_{B},$ (3)
$Z_{A},Z_{B}\sim\mathrm{Zipf}(\alpha),$
where $\pi=\frac{M_{A}}{M_{A}+M_{B}}$ and $M_{A}>M_{B}$.
Based on $Z$, we simulate four types of reasoning trajectories and examine their token-level surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior. More simulation details are provided in § A.6.
Figure 2: Simulation of the student model’s token-level prediction distribution $Z$. Tokens with low probability but relatively high rank—characteristic of the $Z_{B}$ mode—yield smaller token-level rank–surprisal ratios.
[…]
$X_{A}$ (from $Z_{A}$) 0.41 1.38 2.49 1.69
$X_{B}$ (from $Z_{B}$) 0.10 2.73 4.31 1.30
$X_{C}$ (from $Z_{C}$) 0.08 4.73 11.57 2.23
$X_{D}$ (from $Z$) 0.35 1.67 2.93 1.62
Table 3: Simulation results for different types of trajectories, reporting average token probability, surprisal, rank, and $_{\text{token}}. RSR$$_{\text{token}}$ shows promise as a reliable metric for identifying suitable reasoning trajectories (§ 3.4).
Simulation Results
Figure 2 presents the simulated bimodal distribution $Z$, where effective reasoning trajectories are expected to align with the distribution of $Z_{B}$, and thus resemble $X_{B}$. As shown in Table 3, $X_{B}$ exhibits much higher average surprisal (lower absolute probability) than $X_{A}$ and $X_{D}$, while maintaining relatively low rank values (higher-ranked), consistent with the quantitative patterns implied by our analysis. Crucially, the results suggest a promising direction for measuring data suitability, which we formalize in § 3.4.
3.4 Proposed Metric: Rank-Surprisal Ratio
Token-Level Metric
[…]
$\mathrm{RSR}_{\text{token}}(t_{k})\;=\;\frac{\mathrm{Rank}(t_{k})}{\mathrm{Surprisal}(t_{k})}.$ (4)
[…]
From Token-Level to Trajectory-Level
[…]
A natural solution is to exclude tokens with very low surprisal when computing the average. Let $\mathcal{T}_{H}(\mathbf{x})$ denote the set of response tokens whose surprisal lies in the top $H\%$ within a trajectory $\mathbf{x}$. We define a filtered average as
$\mathrm{Avg\text{-}RSR}^{\text{filter}}_{\text{token}}(\mathbf{x})\;=\;\frac{\sum_{t_{k}\in\mathcal{T}_{H}(\mathbf{x})}\mathrm{RSR}_{\text{token}}(t_{k})}{|\mathcal{T}_{H}(\mathbf{x})|}$ (5)
Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2). This suggests that tokens with higher surprisal have a greater impact on student learning and should be emphasized when computing the average.
Accordingly, instead of hard filtering, we adopt a surprisal-weighted average of the token-level ratios. A simple derivation shows that this weighted average is equivalent to a trajectory-level ratio between the sum of token ranks and the sum of token surprisals. For brevity, we denote $r_{k}=\mathrm{Rank}(t_{k})$ and $s_{k}=\mathrm{Surprisal}(t_{k})$:
$\frac{\sum_{k}s_{k}\,\mathrm{RSR}_{\text{token}}(t_{k})}{\sum_{k}s_{k}}=\frac{\sum_{k}s_{k}\,\frac{r_{k}}{s_{k}}}{\sum_{k}s_{k}}\\ =\frac{\sum_{k}\mathrm{Rank}(t_{k})}{\sum_{k}\mathrm{Surprisal}(t_{k})}$ (6)
The resulting metric yields a concise form that can be interpreted as the ratio of average rank to average surprisal at the trajectory level.
In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability. Since tokens with excessively large ranks are effectively indistinguishable from the student’s perspective, we clip rank values at a threshold $r_{max}$. Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR), as
$\mathrm{RSR}(\mathbf{x})\;=\;\frac{\sum_{k}\min\!\bigl(\mathrm{Rank}(t_{k}),r_{max}\bigr)}{\sum_{k}\mathrm{Surprisal}(t_{k})}$ (7)
Interpretation
Our metric admits a simple interpretation. The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior. The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates deviation from dominant patterns and provides informative learning signals. A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
4 Correlation Analysis
Preliminary results in Table 2 have shown that Rank-Surprisal Ratio aligns well with post-training reasoning performance. To provide a more rigorous evaluation and further demonstrate the effectiveness of RSR in measuring data-student suitability, we conduct comprehensive correlation analyses.
4.1 Main Analysis
Metrics (Absolute) Spearman Correlation with Post-Training Performance
Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
Student- Agnostic Teacher Params 0.04 0.34 0.2 0.02 0.26 0.01
Avg-Token Length 0.49 0.68 0.45 0.57 0.47 0.53
Teacher Performance 0.49 0.34 0.13 0.23 0.03 0.23
Verified Accuracy 0.54 0.43 0.25 0.35 0.10 0.33
LLM-judged Quality 0.61 0.52 0.46 0.61 0.40 0.52
Rule-based Quality 0.55 0.56 0.75 0.65 0.75 0.65
Student- Specific Avg-Surprisal 0.24 0.42 0.55 0.55 0.70 0.49
[…]
Avg-Rank 0.41 0.64 0.68 0.61 0.62 0.59
Influence Score 0.52 0.19 0.32 0.47 0.59 0.11
G-Norm 0.44 0.54 0.51 0.57 0.70 0.55
GRACE 0.25 0.58 0.66 0.75 0.69 0.59
Rank-Surprisal Ratio 0.85 0.85 0.92 0.82 0.85 0.86
Table 4: Spearman correlation between different metrics and post-training reasoning performance across student models, reporting absolute values. "Student-Agnostic" metrics are computed independently of the specific student model. Our metric, Rank-Surprisal Ratio, achieves the highest correlation across all students.
For each of the five student models, we compute average trajectory-level suitability (or quality) scores for reasoning datasets generated by eleven teacher models (§ 2.1) using different metrics. We then measure the correlation between these dataset-level scores and the student’s average post-training reasoning performance on the corresponding teacher data. We primarily report Spearman’s correlation coefficient, as Pearson correlation exhibits similar trends. We use a clipping threshold $r_{max}=100$ for our RSR metric in all subsequent experiments. Additional analysis details are provided in § A.7, and complete metric scores are reported in § C.2.
Compared Metrics
We compare RSR against a diverse set of metrics for evaluating reasoning trajectories. These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal; [18]) and rank-based metrics. We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [43], LLM-judged quality scores, and answer accuracy on verifiable questions. In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE; [29]) and influence scores [16].
Results
Table 4 shows that Rank-Surprisal Ratio consistently exhibits strong correlation with post-training reasoning performance across all student models, achieving an average Spearman correlation of 0.86 and outperforming all alternative metrics. These results indicate the effectiveness and practical value of RSR. In contrast, surprisal-based and rank-based metrics alone yield substantially weaker correlations (at most 0.59), highlighting the importance of capturing both informativeness and alignment through the rank-surprisal ratio.
4.2 Ablation Study
Variants Avg. Corr. $\Delta$
Rank-Surprisal Ratio ($r_{max}=100$) $0.856$
   No rank clipping $0.700$ 0.156-0.156
[…]
   Rank clipping: $r_{max}=50$ $0.696$ 0.160-0.160
   Rank clipping: $r_{max}=500$ $0.822$ 0.034-0.034
[…]
   Reduced sample size (200) $0.864$ $0.007$
Table 5: Ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation.
The derivation of Rank-Surprisal Ratio involves several design components as well as a hyperparameter $r_{max}$. We conduct an ablation study to examine how these choices affect correlation performance.
As shown in Table 5, removing either rank clipping or the surprisal-weighted averaging substantially degrades the correlation, validating the necessity of both components in our metric. In addition, the “Reduced sample size” setting estimates the dataset-level RSR using only 200 trajectories per teacher instead of the full 5,000. The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$. More ablation results are provided in § C.3.
5 Practical Applications
Given the reliable data-student suitability estimation provided by Rank-Surprisal Ratio and its strong correlation with post-training performance, we further examine its practical value as a data selection criterion in two representative scenarios.
5.1 Trajectory Selection
Selection Methods Qwen-3-14B L3.1-8B Q2.5-7B Q3-4B Q2.5-3B
AIME24 AIME25 AMC23 MATH500 Avg. Avg. Avg. Avg. Avg.
Random 59.2 46.7 86.2 88.6 70.2 22.1 45.7 53.9 27.9
Token Lengthmax 61.7 51.7 87.5 84.8 71.4 27.3 45.4 51.3 27.1
Rule-based Qualitymax 58.3 47.5 91.3 92.0 72.3 25.8 51.6 58.0 31.2
LLM-judged Qualitymax 60.0 49.2 90.6 93.6 73.4 25.6 51.8 59.1 32.8
Surprisalmin 62.5 50.0 88.1 92.8 73.4 23.5 46.4 53.3 28.9
G-Normmin 59.2 50.0 89.4 92.4 72.7 26.1 49.5 59.1 30.9
Rank-Surprisal Ratiomin 67.5 59.2 93.1 94.6 78.6 28.5 53.2 61.4 34.8
Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods. maxmax and minmin indicate maximizing or minimizing the corresponding metric. Model names are abbreviated as Q for Qwen and L for LLaMA. Additional results, including GPQA evaluation, are provided in § C.4.
Experimental Setting
The trajectory selection task aims to identify the most effective reasoning trajectory from a set of candidates for a given problem or prompt. We adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion. This procedure is repeated for all 5,000 training problems and all student models, yielding student-specific 5,000-trajectory teacher datasets for each selection method. We then fine-tune student models on the constructed datasets and evaluate their post-training reasoning performance on standard math benchmarks, consistent with previous experiments. We compare RSR against a random selection baseline and multiple previously discussed metrics. For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the score.
Results
As shown in Table 6, datasets selected by Rank-Surprisal Ratio consistently achieve the best post-training reasoning performance among all selection methods across student models, demonstrating the effectiveness of RSR in identifying suitable trajectories. Moreover, the performance achieved by RSR is comparable to, and for four students even surpasses, the best performance of any single teacher for each student model (Table 1), which serves as a strong upper bound obtained via brute-force search over teacher datasets. These results underscore the practical value of RSR for selecting effective trajectories prior to training.
Additional trajectory selection results, including complete scores, further analysis, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
5.2 Teacher Selection
Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Avg.
Top-1 Top-2 Top-1 Top-2 Top-1 Top-2 Top-1 Top-2 Top-1 Top-2
Teacher Paramsmax 77.1 71.8 28.1 22.0 47.8 45.0 55.8 53.4 29.6 26.4 45.7
Token Lengthmax 71.8 77.1 22.0 28.1 45.0 47.8 53.4 55.8 26.4 29.6 45.7
Rule-based Qualitymax 72.2 77.1 23.7 28.1 48.2 47.8 56.4 55.8 33.0 29.6 47.2
LLM-judged Qualitymax 71.8 77.2 22.0 26.7 45.0 50.0 53.4 58.8 26.4 31.2 46.3
Surprisalmin 68.8 72.2 22.8 23.7 47.6 48.2 52.2 56.4 30.6 33.0 45.6
GRACEmin 72.2 68.8 22.8 28.1 47.6 48.3 52.2 58.8 30.6 26.4 45.6
Rank-Surprisal Ratiomin 77.2 77.1 26.7 28.1 50.0 47.8 58.8 55.8 31.2 30.6 48.3
Oracle 77.2 77.1 28.1 26.7 50.0 48.2 58.8 56.4 33.0 31.2 48.7
Table 7: Teacher selection results showing post-training reasoning performance of student models trained on data from teachers selected by different methods. Top-1 and Top-2 denote the highest- and second-highest-ranked teachers for each student. "Oracle" corresponds to the ground-truth best and second-best teachers.
Experimental Setting
The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student. We consider a low-resource setting in which generating full training data for every teacher model is infeasible. Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the average trajectory-level score. We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
Results
As shown in Table 7, both the best and second-best teachers selected by Rank-Surprisal Ratio yield strong post-training reasoning performance, achieving average results close to oracle teachers and outperforming other selection methods. Notably, as also observed in our ablation study (Table 5), RSR remains effective when using only 200 trajectories per teacher for measurement, highlighting its robustness and practical value for identifying suitable teachers in low-resource settings.
6 Related Work
Knowledge Distillation
Knowledge distillation is a powerful approach for transferring knowledge from large models to smaller ones [15], and has been widely used in training LLMs [30]. Prior work shows that stronger teachers do not necessarily yield better students, often due to capability gaps [46, 24] or off-policy data [5]. Recent studies address this by better aligning teacher supervision with student behavior, achieving an implicit balance through approaches such as on-policy distillation [2], integration with reinforcement learning [25, 49], SFT optimization [42, 37], teaching assistants [27, 8], and interleaved sampling [38, 31]. By contrast, our work explicitly quantifies the trade-off between informativeness and alignment in distillation, enabling a principled identification of effective teacher data for a given student.
SFT with Reasoning Trajectories
Long CoT trajectories provide strong supervision for improving student models’ reasoning performance via SFT [35, 34]. In line with findings in the general domain that high-quality data improves SFT [51, 41], many studies focus on constructing high-quality CoT data, either by selecting better prompts [44, 40] or by filtering reasoning trajectories [43, 17, 36, 53, 6, 23]. Recognizing that optimal reasoning data may vary across students, recent work explores student-specific trajectory selection strategies [29, 18, 16]. Our work also studies student-specific data selection, but differs from prior work by measuring data–student suitability from the perspective of informative alignment and by conducting more comprehensive evaluations across a wider range of teacher models.
7 Conclusion
In this paper, we study data–student suitability in reasoning distillation and propose Rank-Surprisal Ratio (RSR), a simple metric for identifying suitable reasoning trajectories for a given student. Motivated by our analysis, RSR jointly captures a trajectory’s informativeness and alignment with the student’s behavior, favoring trajectories with low absolute probability but relatively high-ranked tokens. Experiments across diverse teacher–student pairs show that RSR strongly correlates with post-training performance and consistently outperforms existing metrics. We further demonstrate its effectiveness in both trajectory and teacher selection. Overall, our results highlight informative alignment as a promising direction for reasoning distillation.
Limitations
Although our work proposes an effective metric for selecting reasoning trajectories to distill student models, the performance of data selection is inherently constrained by the diversity and quality of candidate trajectories or teacher models. When none of the available teacher trajectories are well suited to a given student, the gains from selection alone may be limited. A promising direction for future work is to use our metric to guide the rewriting or synthesis of reasoning trajectories, rather than selecting from a fixed pool.
In addition, the derived metric takes a simple and intuitive form with a clear interpretation. However, it remains unclear whether it arises from deeper mathematical principles. We have not yet identified a suitable theoretical framework to analyze this aspect, which we leave for future investigation.
Finally, due to resource constraints, we primarily focus on mathematical reasoning tasks with extensive controlled studies, while also including additional evaluation on GPQA. Extending our evaluation to other domains,such as code or additional forms of reasoning, would be valuable. However, since our analysis relies on large-scale trajectory generation and approximately 200 SFT experiments, such extensions would require substantially greater computational resources and are therefore left for future work.
Appendix A Details of Experiments
A.1 Determining Training Problem Set
All teacher trajectory datasets are constructed using a fixed problem set to avoid confounding effects from variations in problem composition. We focus on mathematical reasoning and curate 5,000 math problems from the widely used NuminaMath dataset [21]. Following prior work [20], we apply preprocessing steps such as difficulty filtering to ensure the problem quality.
Specifically, our training dataset consists of problems drawn from the MATH, AIME/AMC, and Olympiads. SkyThought provides a difficulty-labeled version of NuminaMath in which each problem is assigned an integer difficulty score from 0 to 9. Using this scale, we randomly sampled 1,667 problems from the MATH subset with difficulty above three, 1,667 problems from the Olympiads subset with difficulty above eight, and 1,666 problems from the AIME/AMC subset, yielding a balanced training set of 5,000 problems.
We choose 5,000 training problems for three main reasons. First, prior studies have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems. Second, high-quality reasoning problems are relatively scarce, making further scaling less effective in practice. Third, our cross-study incurs quadratic computational cost across teacher–student pairs. Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
A.2 Teacher Trajectory Generation
[…]
If a sampled trajectory exceeds the token budget, we resample up to 10 times; if all attempts still exceed the budget, we truncate the final trajectory. Across all teacher models, truncation rates are below 1%, which helps preserve the quality of training trajectories even for teachers that tend to produce long outputs.
[…]
A.3 Teacher Models and Student Models
Teacher Models
DeepSeek-R1-0528
Qwen-3-235B-Thinking-2507
GPT-OSS-120B (high)
LLaMA-3.3-Nemotron-Super-49B-v1.5
QwQ-32B
Qwen-3-30B-Thinking-2507
Magistral-Small-2506
GPT-OSS-20B (high)
Phi-4-Reasoning-Plus
Qwen-3-8B (thinking)
Qwen-3-4B-Thinking-2507
Table 8: List of teacher models used in our experiments.
As discussed earlier, our experiments adopt a more diverse set of teacher models than prior work. Specifically, we consider the teacher variants shown in Table 8.
For student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B.
A.4 Benchmark Evaluation
We use vLLM together with the Math-Verify package to evaluate post-trained models on mathematics benchmarks. Our evaluated benchmarks, AIME’25, AIME’24, AMC’23, and MATH500, are widely used and span varying difficulty levels, assessing mathematical reasoning and multi-step problem-solving across diverse domains. We additionally conduct evaluation on GPQA-Diamond beyond mathematics, as described in § C.4.2. We adopt the Acc@4 metric as the final score, which averages results over four independent evaluations per problem. Evaluating a single model checkpoint typically takes around half an hour using 8 H200 GPUs.
During inference, we use a temperature of $0.6$, a top_p of $0.95$, and top_k set to 1-1. The maximum generation length is set to 32,768 tokens, consistent with the maximum sequence length used during fine-tuning. Responses that exceed this limit are truncated. This differs from trajectory generation, where we resample multiple times to avoid truncation, as here we aim to evaluate the model’s reasoning ability under a fixed context-length constraint. Under this setting, the comparable performance of the Qwen-3-235B and Qwen-3-30B in Table 1 may be attributable to truncation effects, which we consider a reasonable outcome given the imposed length limit.
A.5 Details of Model Fine-Tuning
Models Learning Rate Batch Size Epoch
Qwen-3-14B 2.0E-05 64 8
LLaMA-3.1-8B 2.0E-05 64 10
Qwen-2.5-7B 2.0E-05 64 10
Qwen-3-4B 2.0E-05 64 10
Qwen-2.5-3B 5.0E-05 64 10
Table 9: Training hyperparameters for different student models.
We use LLaMA-Factory [50] framework to conduct SFT with reasoning trajectory datasets. For different student models, we conduct grid search for best set of hyperparameters. The final setting is reported in the Table 9. During SFT, all student models use a maximum sequence length of 32,768.
Most experiments are conducted on NVIDIA H200 GPUs. A single SFT experiment using the 14B model takes approximately 6 hours on 8 H200 GPUs for training.
For trajectory selection experiments in § 5.1, we fine-tune each selected dataset with three different random seeds and report performance averaged over these runs. For teacher–student distillation experiments in § 2.1, however, each teacher already yields three independent trajectory datasets, so we do not further vary random seeds for fine-tuning.
A.6 Details of Simulation Study
Figure 3: Simulation of the student model’s token-level prediction distribution $Z$, a mixture of Zipf distributions $Z_{A}$ and $Z_{B}$.
In practice, we simulate $Z$ by sampling $M_{A}$ tokens from $Z_{A}$ and $M_{B}$ tokens from $Z_{B}$. We use $M_{A}=1{,}000{,}000$, $M_{B}=250{,}000$, $|\mathcal{V}|=50$, and for each simulated dataset, we sample 10,000 tokens to compute average metrics. Based on a preliminary fit to reasoning data, we set the Zipf exponent to $\alpha=2.3$.
Figure 3 depicts the distributions of $Z_{A}$ and $Z_{B}$ over the vocabulary, with their mixture representing the simulated token-level prediction distribution $Z$ of the student model. Figure 2 is derived from this by ranking the tokens from higher to lower probability.
A.7 Details of Correlation Analysis
We use the Spearman correlation coefficient [45] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance. We also report Pearson correlation results [7] in § C.2.
We report the absolute values of Spearman correlation coefficients in Table 4. For the “Average” values, however, we first average the correlation coefficients across student models and then take the absolute value. We consider this aggregation more appropriate, as correlations with opposite signs across different student models should offset each other; consequently, this averaged value can be lower than the result obtained by averaging absolute correlations.
Since post-training performance is computed by averaging three generation runs per teacher (§ 2.1), we likewise average the dataset-level metric over the three trajectory datasets generated by the same teacher to obtain its final score. We observe that our datset-level metric varies marginally across different datasets generated by the same teacher.
Appendix B Details of Compared Metrics
B.1 LLM-judged Quality
We use Qwen3-32B-Instruct in a Non-Thinking configuration as an automatic judge to evaluate reasoning trajectories. Under a fixed evaluation prompt, the judge produces a structured assessment at two levels. First, it assigns dimension-wise scores over five criteria—Factual Accuracy, Logical Rigor, Solution Completeness, Reasoning Efficiency, and Presentation Quality—each accompanied by a brief justification. Second, it aggregates the dimensional assessments into an overall score in $[0,1]$ together with a concise rationale. For dataset-level comparison, we report the mean of the trajectory-level overall scores. The complete evaluation prompt is given in Table 27, adapted from [9].
B.2 Rule-based Quality
We implement the rule-based criterion used for filtering the LIMO Dataset [43]. Each response is scored using the following weighted indicators: Elaborated reasoning (30%): total word length. Self-Verification (20%): frequency of "check" and "verify". Exploratory Approach (25%): frequency of "perhaps" and "might". Adaptive Granularity (25%): frequency of "therefore" and "since".
To ensure fair comparison across responses of varying lengths, we compute relative keyword frequencies by normalizing absolute keyword counts with respect to the total word count. To account for differences in scale across criteria, we then independently standardize each criterion’s scores into z-scores, which empirically improves correlation.
B.3 Influence Score
We adopt the Influence Score method based on the second-order approximation of Influence Functions [16]. The core principle is to treat the infinitesimal up-weighting of a training sample as a local perturbation to the optimization objective, thereby estimating the induced marginal change in the loss of the evaluation set. In this framework, a higher influence score signifies that the gradient direction of the training sample aligns with minimizing the evaluation loss in the neighborhood of the converged parameters.
We implement this baseline using the Kronfluence framework. To make the computation tractable for LLMs, we employ the EK-FAC strategy to approximate the Hessian matrix [10]. Specifically, we first precompute the EK-FAC factors on a reference model. Since we do not know which trajectory is better a priori, we use the model trained with randomly selected trajectories per problem as the reference model, corresponding to the “Random” variant in Table 6. Then, we compute the pairwise influence scores between the training samples and the evaluation dataset. To optimize for memory and computational efficiency, we apply a low-rank approximation to the query gradients (rank =4=4) and utilize bfloat16 precision for the Inverse Hessian-Vector Product calculations. Finally, we average these pairwise scores across the evaluation set to obtain the sample-level influence score $s(d)$, which serves as our baseline for ranking data suitability.
B.4 G-Norm
G-Norm [29] assesses the suitability of generated data for a student model by measuring the magnitude of the student’s local gradient signal induced by the data. Formally, G-Norm calculates the magnitude of the loss gradient with respect to the student model’s parameters. To address the computational constraints associated with high-dimensional parameter spaces, the method utilizes random projection for dimensionality reduction. For each reasoning trajectory $(x,y)$, we compute the gradient of the loss derived from the student model and apply length normalization to eliminate bias towards longer sequences. These gradients are then projected onto a lower-dimensional subspace via a fixed random matrix, followed by the computation of their $L_{2}$ norms.
B.5 GRACE
We adopt GRACE [29] as a baseline metric for further evaluating the generated reasoning trajectory dataset from a gradient-based perspective. GRACE characterizes the geometry of the optimization landscape by analyzing the spectral structure of the student model gradients. This approach allows for a holistic assessment of how effectively the generated data spans the parameter space required for model optimization.
The calculation proceeds by first projecting the high-dimensional gradients onto a lower-dimensional subspace via a fixed random matrix to ensure computational feasibility. To address the empirical tendency of gradient norms to diminish in longer sequences, the method rescales the projected vectors logarithmically based on the response length. GRACE then computes the metric using a cross-validation strategy where the dataset is divided into multiple partitions. The gradients from a held-out partition are weighted by the regularized inverse covariance matrix estimated from the remaining data. This process effectively quantifies the expected squared norm of the gradients after whitening them with the estimated spectrum of the distribution.
Because GRACE’s cross-validation strategy yields a dataset-level metric rather than a per-sample score, we apply GRACE for teacher selection but not for trajectory-level selection.
B.6 Others
Here we give formal definitions for Avg-Surprisal and Avg-Surplocal. Given a trajectory $x=(t_{1},\ldots,t_{T})$,
$\mathrm{Avg\_Surprisal}(\mathbf{x})=\frac{1}{T}\sum_{k=1}^{T}-\log p_{\theta}(t_{k}\mid\mathbf{c}_{k})$ (8)
$\mathrm{Avg\_Surp_{local}}(\mathbf{x})=\frac{1}{T}\sum_{k=1}^{T}-\log p_{\theta}(t_{k}\mid\mathbf{c}_{k}^{\text{local}})$ (9)
Appendix C More Results and Analysis
C.1 Computational Cost for Rank-Surprisal Ratio
The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model. During the forward pass, we collect each token’s surprisal and rank from the model logits. Computing Rank-Surprisal Ratio over the 5,000-trajectory dataset typically takes about one hour on a single H200 GPU, which is significantly cheaper than SFT.
C.2 Additional Results for Correlation Analysis
Metrics Student Models
Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
Rank-Surprisal Ratio Spearman 0.855 0.845 0.918 0.818 0.845 0.856
Pearson 0.654 0.880 0.805 0.819 0.811 0.794
Table 10: Comparison of Spearman and Pearson correlation results on Rank-Surprisal Ratio (absolute values).
Table 10 compares Spearman and Pearson correlations. The results show that Rank-Surprisal Ratio also exhibits strong Pearson correlation with post-training performance.
Table 22, 23, 24, 25, and 26 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
C.3 Additional Ablation Study
Variants Avg. Corr. $\Delta$
Rank-Surprisal Ratio $0.856$
   Fixed student $0.785$ 0.071-0.071
   use Rank1.05 $0.845$ 0.011-0.011
   use Rank0.95 $0.787$ 0.069-0.069
   use Surprisal1.05 $0.847$ 0.009-0.009
   use Surprisal0.95 $0.873$ $0.017$
Avg-Rank (clip at $r_{max}=100$) $0.552$ 0.304-0.304
Rank Minus Surprisal $0.585$ 0.271-0.271
Rank-Entropy Ratio $0.764$ 0.092-0.092
Table 11: Additional ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation.
We conduct some additional ablation studies. Results are shown in Table 11.
The “Fixed student” variant computes RSR using a fixed model (Qwen-3-14B) instead of the target student, and the resulting drop in correlation highlights the importance of student-specific estimation. “Avg-Rank (clip at $r_{max}=100$)” applies rank clipping to Avg-Rank, but it still fails to achieve better correlation than rank alone. “Rank Minus Surprisal” is computed by subtracting the surprisal from the (clipped) rank. Although this metric also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance. Rank–Entropy Ratio is computed as the ratio between the (clipped) average rank and the average token entropy, where entropy is loosely related to surprisal. While it achieves good correlation, its correlation is still lower than that of RSR.
We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to power 1 for both rank and surprisal), for example "use Rank1.05". The results indicate that our metric is robust to these variations and may yield higher correlation with hyperparameter tuning. Nevertheless, we use the default power-1 setting to keep the formulation simple.
C.4 Additional Results for Trajectory Selection
C.4.1 Ablation Study on Trajectory Selection
Variant Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
Math Avg. Math Avg. Math Avg. Math Avg. Math Avg.
Rank-Surprisal Ratiomin 78.6 28.5 53.2 61.4 34.8 51.3
   With Correctness Filtering 77.5 27.9 52.3 60.7 34.9 50.7
   Fewer Candidates per Teacher 77.6 27.8 52.6 60.9 33.8 50.5
Table 12: Ablation study of trajectory selection with RSRmin-based variant methods across student models. "Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
Table 12 presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
For "With Correctness Filtering", we consider a combined setting that uses both RSR and verified correctness for trajectory selection. Specifically, for problems with verifiably correct trajectories, we first discard incorrect ones and then select among the remaining correct trajectories based on RSR. The results show no significant improvement over selecting trajectories solely based on RSR, particularly for larger student models. This suggests that correctness is sometimes less critical than overall data suitability.
For "Fewer Candidates per Teacher", we evaluate an 11-to-1 setting in which each candidate pool contains 11 teacher trajectories (one per teacher), instead of the original 33-to-1 setting (three per teacher). This setting focuses on selecting the best trajectory for each problem across different teachers, rather than across multiple generations from the same teacher. The results are comparable, with a slight performance gap relative to the 33-to-1 setting, indicating that RSR remains effective even when each teacher provides only a single trajectory. These findings further suggest that while RSR captures suitability differences across generations from the same teacher, such differences are less pronounced than those across different teacher models.
C.4.2 Additional Evaluation on GPQA and Full Results
Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
GPQA GPQA GPQA GPQA GPQA
Random 48.5 26.8 35.4 43.4 22.2 35.3
LLM-judged Qualitymax 53.5 27.3 35.4 45.5 21.2 36.6
Rank-Surprisal Ratiomin 55.1 31.3 38.9 45.5 31.3 40.4
Table 13: Comparison of different trajectory selection methods on the GPQA-Diamond benchmark across student models.
To more comprehensively evaluate the impact of different trajectory selection methods on post-trained models’ reasoning capabilities beyond mathematical problems, we conduct additional evaluation on the GPQA-Diamond benchmark [33]. GPQA-Diamond consists of 198 challenging multiple-choice questions spanning biology, physics, and chemistry.
Table 13 summarizes the evaluation results of different trajectory selection methods on GPQA-Diamond. Although the results are less stable than those on mathematical benchmarks due to the out-of-domain nature of this evaluation, datasets selected by RSR still achieve the best overall post-training performance. This suggests that RSR can identify suitable trajectories that consistently improve student models’ general reasoning capabilities, even when training solely on mathematical problems.
The complete trajectory selection results underlying Table 6 are presented in Table 17, 18, 19, 20, and 21.
C.4.3 Analysis of Datasets Selected by RSR
Selected Datasets Data Composition over Teacher Models Metrics
Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR
RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57
RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69
RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67
RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56
RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73
Table 14: Data composition and metrics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) across different student models. Model names are abbreviated as Q for Qwen and L for LLaMA.
Table 14 shows the data composition of datasets selected by RSR across 11 teacher models. The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students. QwQ-32B is generally preferred across student models, consistent with its stable performance. For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed.
The RSR values of the selected datasets are also reported in Table 14. These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § C.2 for complete metrics), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
Selection Methods Qwen-3-14B Qwen-2.5-7B
Math Avg. Math Avg.
Random 72.8 47.3
LLM-judged Qualitymax 74.4 48.3
Rank-Surprisal Ratiomin 76.8 50.0
Table 15: Comparison of different trajectory selection methods under a reduced-teacher setting.
To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers. Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B. This teacher set is formed by combining the teachers used in Table 2 and 7.
The results are shown in Table 15. Datasets selected by RSR still achieve superior post-training reasoning performance compared with the baselines, demonstrating the effectiveness of our metric when only a limited number of teachers are available. We also observe that the performance gap narrows when selecting from seven teachers compared with eleven teachers, which is expected and suggests that a larger candidate space enables trajectory selection to more effectively identify high-quality training data.
Moreover, Table 16 presents the data composition of datasets selected by RSR from the reduced set of seven teacher models. The distributions differ markedly for Qwen-3-14B and Qwen-2.5-7B: the former student model tends to favor teachers such as DeepSeek-R1 and Qwen-3-30B, whereas the latter student model shows a stronger preference for smaller models such as Qwen-3-8B. These results further demonstrate the effectiveness of RSR in selecting teacher trajectories that are well suited to specific student models.
Selected Datasets Data Composition over Teacher Models
Deepseek-R1 Qwen-3-235B Nemotron Qwen-3-30B Magistral GPT-OSS-20B Qwen-3-8B
RSRmin on Qwen-3-14B 27.40% 21.00% 1.42% 28.84% 4.92% 0.24% 16.18%
RSRmin on Qwen-2.5-7B 14.52% 8.60% 6.94% 20.76% 10.64% 0.14% 38.40%
Table 16: Data composition of trajectory datasets selected by RSR under a reduced-teacher setting (see the setting in § C.4.4 and results in Table 15).
Appendix D Others
D.1 License for Artifacts and Data Consent
All artifacts used in this paper are publicly available for academic research purposes, including AIME, AMC, MATH500, and NuminaMath.
D.2 Data Statement
The training datasets consist solely of mathematics problems and solutions and contain no offensive content or personal information.
D.3 AI Assistant Usage Statement
We used ChatGPT for writing refinement and minor coding assistance. AI assistants were not involved in research innovation, and all core contributions were developed solely by the authors.
Selection Methods Qwen-3-14B
AIME’24 AIME’25 AMC’23 MATH500 Math Avg. GPQA-Diamond
Random 59.2 46.7 86.3 88.6 70.2 48.5
Token Lengthmax 61.7 51.7 87.5 84.8 71.4
Rule-based Qualitymax 58.3 47.5 91.3 92.0 72.3
LLM-judged Qualitymax 60.0 49.2 90.6 93.6 73.4 53.5
Surprisalmin 62.5 50.0 88.1 92.8 73.4
G-Normmin 59.2 50.0 89.4 92.4 72.7
Rank-Surprisal Ratiomin 67.5 59.2 93.1 94.6 78.6 55.1
Table 17: Full post-training evaluation results for trajectory selection on Qwen-3-14B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
Selection Methods LLaMA-3.1-8B
AIME’24 AIME’25 AMC’23 MATH500 Math Avg. GPQA-Diamond
Random 2.5 5.8 29.4 50.8 22.1 26.8
Token Lengthmax 8.3 6.7 36.9 57.4 27.3
Rule-based Qualitymax 6.7 9.2 29.4 58.0 25.8
LLM-judged Qualitymax 1.7 4.2 38.1 58.6 25.6 27.3
Surprisalmin 5.0 6.7 25.6 56.8 23.5
G-Normmin 5.8 4.2 36.9 57.6 26.1
Rank-Surprisal Ratiomin 5.0 8.3 36.9 63.6 28.5 31.3
Table 18: Full post-training evaluation results for trajectory selection on LLaMA-3.1-8B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
Selection Methods Qwen-2.5-7B
AIME’24 AIME’25 AMC’23 MATH500 Math Avg. GPQA-Diamond
Random 18.3 19.2 62.5 82.8 45.7 35.4
Token Lengthmax 22.5 21.7 61.9 75.6 45.4
Rule-based Qualitymax 24.2 25.0 72.5 84.8 51.6
LLM-judged Qualitymax 30.0 23.3 66.9 87.0 51.8 35.4
Surprisalmin 22.5 20.8 63.1 79.2 46.4
G-Normmin 27.5 25.0 66.3 79.2 49.5
Rank-Surprisal Ratiomin 29.2 25.8 71.3 86.6 53.2 38.9
Table 19: Full post-training evaluation results for trajectory selection on Qwen-2.5-7B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
Selection Methods Qwen-3-4B
AIME’24 AIME’25 AMC’23 MATH500 Math Avg. GPQA-Diamond
Random 30.8 30.8 68.8 85.2 53.9 43.4
Token Lengthmax 28.3 28.3 71.9 76.6 51.3
Rule-based Qualitymax 36.7 32.5 77.5 85.4 58.0
LLM-judged Qualitymax 37.5 32.5 77.5 88.8 59.1 45.5
Surprisalmin 29.2 33.3 71.9 78.8 53.3
G-Normmin 34.2 33.3 79.4 89.4 59.1
Rank-Surprisal Ratiomin 44.2 35.0 77.5 88.8 61.4 45.5
Table 20: Full post-training evaluation results for trajectory selection on Qwen-3-4B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
Selection Methods Qwen-2.5-3B
AIME’24 AIME’25 AMC’23 MATH500 Math Avg. GPQA-Diamond
Random 6.7 7.5 36.3 61.2 27.9 22.2
Token Lengthmax 5.0 10.0 37.5 56.0 27.1
Rule-based Qualitymax 5.8 10.0 45.0 63.8 31.2
LLM-judged Qualitymax 12.5 10.8 39.4 68.4 32.8 21.2
Surprisalmin 5.8 8.3 39.4 62.0 28.9
G-Normmin 9.2 9.2 46.3 59.0 30.9
Rank-Surprisal Ratiomin 11.7 11.7 45.0 70.8 34.8 31.3
Table 21: Full post-training evaluation results for trajectory selection on Qwen-2.5-3B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
Metrics Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ-32B Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B
Avg-Token Length 12077.7 12571.1 2552.3 8798.2 9070.4 10887.1 10993.3 3822.7 3643.1 10473.9 13145.8
Teacher Performance 0.911 0.912 0.883 0.823 0.852 0.923 0.710 0.834 0.727 0.825 0.873
Verified Accuracy 0.849 0.859 0.801 0.786 0.820 0.857 0.776 0.784 0.804 0.803 0.835
Rule-based Quality 0.226 -0.031 -0.484 0.259 0.395 -0.106 0.100 -0.393 -0.389 0.465 -0.042
LLM-judged Quality 0.908 0.966 0.896 0.882 0.901 0.963 0.815 0.863 0.823 0.911 0.951
G-Norm 33.615 33.581 72.106 31.801 39.701 34.569 35.455 66.665 62.448 33.913 33.580
Influence Score ($\times 10^{5}$) 0.161 0.696 0.418 1.447 0.795 0.671 -0.008 0.244 -0.348 0.816 0.633
Avg-Surprisal 0.660 0.616 1.162 0.424 0.629 0.591 0.410 1.276 1.068 0.485 0.581
Avg-Rank 49.412 55.609 430.379 65.293 55.249 58.448 41.968 365.886 303.840 45.098 49.145
GRACE 0.028 0.028 0.168 0.025 0.031 0.030 0.025 0.154 0.113 0.023 0.026
[…]
RSR (200 sample) 2.916 2.943 3.504 3.342 2.684 2.915 3.252 3.679 3.348 2.961 2.904
RSR 2.925 2.940 3.527 3.352 2.673 2.923 3.302 3.645 3.360 3.003 2.918
Table 22: Full metric assessment results on Qwen-3-14B. Model names are abbreviated with Q for Qwen.
Metrics Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ-32B Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B
Avg-Token Length 12077.7 12571.1 2552.3 8798.2 9070.4 10887.1 10993.3 3822.7 3643.1 10473.9 13145.8
Teacher Performance 0.911 0.912 0.883 0.823 0.852 0.923 0.710 0.834 0.727 0.825 0.873
Verified Accuracy 0.849 0.859 0.801 0.786 0.820 0.857 0.776 0.784 0.804 0.803 0.835
Rule-based Quality 0.226 -0.031 -0.484 0.259 0.395 -0.106 0.100 -0.393 -0.389 0.465 -0.042
LLM-judged Quality 0.908 0.966 0.896 0.882 0.901 0.963 0.815 0.863 0.823 0.911 0.951
G-Norm 52.768 55.501 120.459 55.882 56.478 57.108 43.666 109.119 95.909 48.365 53.058
Influence Score ($\times 10^{6}$) 2.865 1.330 2.361 1.123 1.359 1.258 0.932 2.755 2.205 1.125 1.243
Avg-Surprisal 0.945 0.927 1.418 0.724 0.953 0.899 0.668 1.530 1.277 0.754 0.866
Avg-Rank 10.328 11.101 72.356 11.762 11.028 11.206 8.463 66.664 49.643 8.721 9.757
GRACE 0.162 0.177 1.005 0.185 0.183 0.181 0.111 0.853 0.639 0.143 0.159
[…]
RSR (200 sample) 2.995 3.044 3.976 2.997 2.848 2.960 3.000 4.104 3.608 2.857 2.946
RSR 2.996 3.044 3.971 3.016 2.818 2.965 3.020 4.038 3.633 2.882 2.945
Table 23: Full metric assessment results on LLaMA-3.1-8B. Model names are abbreviated with Q for Qwen.
Metrics Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ-32B Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B
Avg-Token Length 12077.7 12571.1 2552.3 8798.2 9070.4 10887.1 10993.3 3822.7 3643.1 10473.9 13145.8
Teacher Performance 0.911 0.912 0.883 0.823 0.852 0.923 0.710 0.834 0.727 0.825 0.873
Verified Accuracy 0.849 0.859 0.801 0.786 0.820 0.857 0.776 0.784 0.804 0.803 0.835
Rule-based Quality 0.226 -0.031 -0.484 0.259 0.395 -0.106 0.100 -0.393 -0.389 0.465 -0.042
LLM-judged Quality 0.908 0.966 0.896 0.882 0.901 0.963 0.815 0.863 0.823 0.911 0.951
G-Norm 42.327 39.127 84.024 38.592 45.587 40.408 32.527 78.025 73.562 38.400 37.860
Influence Score ($\times 10^{6}$) -0.520 -0.766 -0.812 -0.732 -0.737 -0.767 -0.466 -0.788 -0.740 -0.727 -0.770
Avg-Surprisal 0.825 0.799 1.236 0.597 0.820 0.767 0.553 1.356 1.131 0.647 0.748
Avg-Rank 6.192 6.438 36.628 6.413 6.312 6.411 4.724 35.233 25.818 5.000 5.683
GRACE 0.168 0.199 1.760 0.126 0.160 0.142 0.119 1.491 0.908 0.120 0.139
[…]
RSR (200 sample) 2.999 3.030 3.670 3.066 2.803 2.950 3.067 3.887 3.471 2.864 2.935
RSR 3.002 3.023 3.686 3.086 2.779 2.951 3.091 3.827 3.468 2.888 2.940
Table 24: Full metric assessment results on Qwen-2.5-7B. Model names are abbreviated with Q for Qwen.
Metrics Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ-32B Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B
Avg-Token Length 12077.7 12571.1 2552.3 8798.2 9070.4 10887.1 10993.3 3822.7 3643.1 10473.9 13145.8
Teacher Performance 0.911 0.912 0.883 0.823 0.852 0.923 0.710 0.834 0.727 0.825 0.873
Verified Accuracy 0.849 0.859 0.801 0.786 0.820 0.857 0.776 0.784 0.804 0.803 0.835
Rule-based Quality 0.226 -0.031 -0.484 0.259 0.395 -0.106 0.100 -0.393 -0.389 0.465 -0.042
LLM-judged Quality 0.908 0.966 0.896 0.882 0.901 0.963 0.815 0.863 0.823 0.911 0.951
G-Norm 33.263 33.036 70.922 35.749 43.248 33.884 39.333 67.150 62.701 38.211 33.109
Influence Score ($\times 10^{4}$) -0.284 -0.013 -1.077 3.601 1.919 0.068 -1.214 -1.202 -0.965 1.395 -1.149
Avg-Surprisal 0.721 0.685 1.208 0.474 0.693 0.650 0.450 1.319 1.102 0.526 0.615
Avg-Rank 9.172 10.267 82.658 10.747 9.773 10.346 7.503 72.151 54.562 7.710 9.011
GRACE 0.148 0.151 0.524 0.148 0.165 0.145 0.139 0.436 0.425 0.124 0.116
[…]
RSR (200 sample) 2.947 2.961 3.771 3.197 2.660 2.908 3.124 3.937 3.579 2.881 2.919
RSR 2.958 2.955 3.794 3.216 2.649 2.917 3.160 3.888 3.588 2.919 2.928
Table 25: Full metric assessment results on Qwen-3-4B. Model names are abbreviated with Q for Qwen.
Metrics Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ-32B Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B
Avg-Token Length 12077.7 12571.1 2552.3 8798.2 9070.4 10887.1 10993.3 3822.7 3643.1 10473.9 13145.8
Teacher Performance 0.911 0.912 0.883 0.823 0.852 0.923 0.710 0.834 0.727 0.825 0.873
Verified Accuracy 0.849 0.859 0.801 0.786 0.820 0.857 0.776 0.784 0.804 0.803 0.835
Rule-based Quality 0.226 -0.031 -0.484 0.259 0.395 -0.106 0.100 -0.393 -0.389 0.465 -0.042
LLM-judged Quality 0.908 0.966 0.896 0.882 0.901 0.963 0.815 0.863 0.823 0.911 0.951
G-Norm 29.027 27.225 74.196 27.733 30.065 27.843 26.490 67.096 62.386 25.996 26.518
Influence Score ($\times 10^{5}$) -1.288 -2.927 -2.157 -2.873 -3.711 -2.730 -1.694 -2.311 -2.555 -3.225 -2.812
Avg-Surprisal 0.903 0.885 1.346 0.657 0.891 0.847 0.608 1.454 1.210 0.704 0.822
Avg-Rank 18.247 19.034 145.459 24.048 21.422 19.906 16.483 119.646 90.940 16.967 17.978
GRACE 0.096 0.092 0.602 0.093 0.100 0.147 0.077 0.991 0.426 0.075 0.085
[…]
RSR (200 sample) 3.094 3.109 3.929 3.084 2.885 3.017 3.029 4.095 3.603 2.891 2.991
RSR 3.095 3.103 3.944 3.107 2.860 3.012 3.050 4.037 3.622 2.911 2.994
Table 26: Full metric assessment results on Qwen-2.5-3B. Model names are abbreviated with Q for Qwen.
Prompt
Table 27: Evaluation prompt for LLM-judged quality assessment.

### v2

@@
 Long chain-of-thought (CoT) trajectories provide rich supervision signals for distilling reasoning from teacher to student LLMs.
-However, both prior work and our experiments show that trajectories from stronger teachers do not necessarily yield better students, highlighting the importance of data–student suitability in distillation.
-Existing methods assess suitability primarily through student likelihood, favoring trajectories that closely align with the model’s current behavior but overlooking more informative ones.
+However, both prior work and our experiments show that trajectories from stronger teachers do not necessarily yield better students, highlighting the importance of data-student suitability in distillation.
+Existing methods assess suitability primarily through student likelihood, favoring trajectories that align closely with the student model’s current behavior but overlooking more informative ones.
 Addressing this, we propose Rank-Surprisal Ratio (RSR), a simple metric that captures both alignment and informativeness to assess the suitability of a reasoning trajectory.
-RSR is motivated by the observation that effective trajectories typically combine low absolute probability with relatively high-ranked tokens under the student model, balancing learning signal strength and behavioral alignment.
+RSR is motivated by the observation that effective trajectories typically balance learning signal strength and behavioral alignment by combining low absolute probability with relatively high-ranked tokens under the student model.
 Concretely, RSR is defined as the ratio of a trajectory’s average token-wise rank to its average negative log-likelihood, and is straightforward to compute and interpret.
-Across five student models and reasoning trajectories from 11 diverse teachers, RSR strongly correlates with post-training performance (average Spearman 0.86), outperforming existing metrics.
+Across five student models and reasoning trajectories from 11 diverse teachers, RSR strongly correlates with post-training reasoning performance (average Spearman 0.86), consistently outperforming existing metrics.
 We further demonstrate its practical utility in both trajectory selection and teacher selection.
-$\dagger$ Corresponding authors. Inquiries may be sent to: yumingyang23@m.fudan.edu.cn, zouyicheng@pjlab.org.cn, qz@fudan.edu.cn* Code is available at https://github.com/UmeanNever/RankSurprisalRatio.
+$\dagger$ Corresponding authors.
+Inquiries may be sent to: yumingyang23@m.fudan.edu.cn, zouyicheng@pjlab.org.cn, qz@fudan.edu.cn* Code and data are available at https://github.com/UmeanNever/RankSurprisalRatio.
 1 Introduction
@@
 In contrast, more informative trajectories are typically less familiar to the student and thus overlooked by these methods.
+To facilitate more effective learning, it is crucial to strike a balance between familiarity and informativeness, echoing the psychological concept of the zone of proximal development [36].
 This leads to a fundamental Informative Alignment challenge: how to identify reasoning data that are both well aligned with the student and sufficiently informative?
 To address this challenge, we propose a simple yet effective metric, Rank-Surprisal Ratio (RSR), which quantifies the suitability of a reasoning trajectory for a given student by jointly capturing alignment and informativeness.
-Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting absolute unfamiliarity and relative familiarity.
+Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting both absolute unfamiliarity and relative familiarity.
 Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while still containing behavioral patterns within the student’s prior experience, such that their tokens rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
 Figure 1: Illustration of the intuition behind RSR.
-Suitable reasoning trajectories should balance informativeness and alignment by having low absolute probability but relatively high-ranked tokens under the student model.
+Suitable reasoning trajectories should balance informativeness and alignment, having low absolute probability while their tokens remain relatively high-ranked under the student model.
 Based on this insight and consistent numerical patterns observed in simulation studies, we define our suitability metric, Rank-Surprisal Ratio, as the ratio between a trajectory’s average token-wise rank [^1] Higher-ranked tokens have lower rank values.
@@
 Our main contributions are three-fold:
-  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data–student suitability.
-  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
+  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data-student suitability.
+  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory for a given student model by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
   • We demonstrate the practical utility of Rank-Surprisal Ratio in two data engineering scenarios, trajectory selection and teacher selection, where it serves as an effective criterion and outperforms existing methods.
@@
 (2) For each teacher–student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance.
-All students are pretrained base models.
+All students are pre-trained base models.
 To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher–student pair.
 Reported results are averaged over these three runs.
-More implementation details are provided in the appendix.
+More implementation details are provided in Appendix A.
 Teachers
-We use 11 reasoning LLMs spanning 4B to 671B parameters across multiple model families, including GPT-OSS [3], DeepSeek [12], Qwen [39], LLaMA-Nemotron [4], and Phi [1].
+We use 11 reasoning LLMs (§ A.3) spanning 4B to 671B parameters across multiple model families, including DeepSeek [13], GPT-OSS [3], Qwen [41], LLaMA-Nemotron [4], and Phi [1].
 Benchmarks
-We evaluate the reasoning performance of the fine-tuned student models on four standard math benchmarks: AIME’25, AIME’24, AMC’23, and MATH500 [14] using the Acc@4 metric, and report results averaged across all benchmarks.
+We evaluate the reasoning performance of fine-tuned student models on four popular math benchmarks (AIME’25, AIME’24, AMC’23, and MATH500 [15]) using the Acc@4 metric (§ A.4), and report results averaged across all benchmarks.
 2.2 Results
@@
 The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model.
-Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) may likewise be ineffective at improving stronger students (Qwen-3-14B).
+Pairing strong teachers (Deepseek-R1) with much weaker students (Qwen-2.5-3B) often fails to yield strong performance, while weaker teachers (Qwen-3-4B-Thinking) can likewise be ineffective when training stronger students (Qwen-3-14B).
 Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb.
@@
 3.1 Surprisal and Rank
-We introduce two different methods to quantify the amount of information carried by a token with respect to a student model. They serve as building blocks of our metric.
+We first introduce two fundamental methods for quantifying the amount of information a token carries with respect to a student model. They serve as the building blocks of our metric.
 Surprisal (Negative Log-Likelihood)
 A common measure is based on the probability of generating the current token $t_{k}$ given its preceding context $\mathbf{c}_{k}=(t_{1},\ldots,t_{k-1})$ under the student model $\theta$.
-For numerical stability, probabilities are typically transformed into the log space, and the negative log-likelihood is used as a measure of informativeness, also known as Surprisal [13]
+For numerical stability, probabilities are typically transformed into log space, and the negative log-likelihood—also known as surprisal—is used as a measure of informativeness [14].
 $\mathrm{Surprisal}(t_{k})=-\log p_{\theta}\bigl(t_{k}\mid\mathbf{c}_{k}\bigr)$ (1)
@@
 Unlike surprisal, rank captures relative familiarity of the token by comparing target token against alternative candidates, revealing signals overlooked by probability-based measures.
-For instance, a token may receive low absolute probability while still being among the top-ranked candidates.
+For instance, a token may be assigned a low absolute probability while still ranking among the top candidates.
 3.2 Limitations of Probability-Based Metrics
-Existing work primarily relies on surprisal or log-probability to assess data suitability.
+Existing work primarily relies on log-probability or surprisal to assess data suitability.
 For example, Zhang et al.
-[47] selects trajectories based on the average log-probability of response tokens under the student model.
+[49] select trajectories based on the average log-probability of response tokens under the student model.
 Since surprisal is the negation of log-probability, we implement this metric as the average surprisal, denoted as Avg-Surprisal.
 More recently, Just et al.
-[18] computes token-level log-probability based on a local context $\mathbf{c}_{k}^{\text{local}}$ (several preceding sentences), which we implement as average local surprisal (Avg-Surplocal).
-Under these metrics, trajectories with lower surprisal are considered more suitable.
+[19] compute token-level log-probability based on a local context $\mathbf{c}_{k}^{\text{local}}$ (several preceding sentences), which we implement as average local surprisal (Avg-Surplocal, § B.6).
+Under these metrics, trajectories with lower surprisal are considered more suitable for the student model.
 However, as shown in Table 2, lower surprisal (i.e., higher probability) does not necessarily lead to better post-training reasoning performance.
@@
 3.3 Insight and Simulation
-The above analysis suggests that effective teacher trajectories should balance the informativeness of the data with their alignment to the student’s current behavior: they should be neither overly similar to the student’s own generations nor excessively deviant from its predicted distribution.
+The above analysis suggests that suitable (i.e., effective) teacher trajectories should strike a balance between data informativeness and alignment with the student’s current behavior: they should be neither overly similar to the student’s own generations nor excessively deviant from its prediction distribution.
 At first glance, this balance may appear to pose a dilemma.
@@
 Informativeness does not require trajectories to be entirely unfamiliar; rather, it suffices that they deviate from the dominant patterns and thus have low absolute probability of being generated by the student.
-Conversely, alignment does not require exact agreement with the student’s outputs, but instead that the corresponding tokens have relatively higher likelihoods than other candidates in the vocabulary.
+Conversely, alignment does not entail exact agreement with the student’s outputs, but instead requires that the corresponding tokens have relatively higher likelihoods than other candidates in the vocabulary.
 Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while still exhibiting behavioral patterns within the student’s prior experience.
-Therefore, tokens in such trajectories are assigned low absolute probability (high surprisal) by the student model while still ranking relatively high (low rank values) in its prediction distribution.
+Therefore, tokens in such trajectories are assigned low absolute probability (i.e., high surprisal) by the student model while still ranking relatively high (i.e., having low rank values) in its prediction distribution.
 To validate these quantitative patterns and identify features that characterize effective learning trajectories, we conduct a simulation study.
 Simulation Setting
-We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we consider effective.
+We simulate the student model’s token-level prediction distribution for reasoning trajectories and examine the numerical patterns exhibited by teacher trajectories that we deem effective.
 Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$.
@@
 where $\pi=\frac{M_{A}}{M_{A}+M_{B}}$ and $M_{A}>M_{B}$.
-Based on $Z$, we simulate four types of reasoning trajectories and examine their token-level surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior.
+Based on $Z$, we simulate four types of reasoning trajectories and examine their average token surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior.
 More simulation details are provided in § A.6.
@@
 $X_{D}$ (from $Z$) 0.35 1.67 2.93 1.62
-Table 3: Simulation results for different types of trajectories, reporting average token probability, surprisal, rank, and $_{\text{token}}.
-RSR$$_{\text{token}}$ shows promise as a reliable metric for identifying suitable reasoning trajectories (§ 3.4).
+[…]
 Simulation Results
 Figure 2 presents the simulated bimodal distribution $Z$, where effective reasoning trajectories are expected to align with the distribution of $Z_{B}$, and thus resemble $X_{B}$.
-As shown in Table 3, $X_{B}$ exhibits much higher average surprisal (lower absolute probability) than $X_{A}$ and $X_{D}$, while maintaining relatively low rank values (higher-ranked), consistent with the quantitative patterns implied by our analysis.
+As shown in Table 3, $X_{B}$ exhibits a much higher average surprisal (lower probability) than $X_{A}$ and $X_{D}$, while maintaining relatively low rank values (top-ranked), consistent with the quantitative patterns implied by our prior analysis.
 Crucially, the results suggest a promising direction for measuring data suitability, which we formalize in § 3.4.
@@
 Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2).
-This suggests that tokens with higher surprisal have a greater impact on student learning and should be emphasized when computing the average.
+This suggests that tokens with higher surprisal have a greater impact on student learning and should therefore be emphasized when computing the average.
 Accordingly, instead of hard filtering, we adopt a surprisal-weighted average of the token-level ratios.
@@
 $\frac{\sum_{k}s_{k}\,\mathrm{RSR}_{\text{token}}(t_{k})}{\sum_{k}s_{k}}=\frac{\sum_{k}s_{k}\,\frac{r_{k}}{s_{k}}}{\sum_{k}s_{k}}\\ =\frac{\sum_{k}\mathrm{Rank}(t_{k})}{\sum_{k}\mathrm{Surprisal}(t_{k})}$ (6)
-The resulting metric yields a concise form that can be interpreted as the ratio of average rank to average surprisal at the trajectory level.
+The resulting metric yields a concise form that can be interpreted as the ratio of the average token rank to the average token surprisal over a trajectory.
 In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability.
 Since tokens with excessively large ranks are effectively indistinguishable from the student’s perspective, we clip rank values at a threshold $r_{max}$.
-Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR), as
+Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR [^2] By default, RSR refers to the trajectory-level RSR in this paper, unless otherwise specified (e.g., in correlation analysis).), as
 $\mathrm{RSR}(\mathbf{x})\;=\;\frac{\sum_{k}\min\!\bigl(\mathrm{Rank}(t_{k}),r_{max}\bigr)}{\sum_{k}\mathrm{Surprisal}(t_{k})}$ (7)
@@
 The numerator, Rank, captures relative familiarity: lower rank values indicate that tokens in this trajectory are preferred among alternative candidates by the student and align with the model’s existing behavior.
-The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates deviation from dominant patterns and provides informative learning signals.
+The denominator, Surprisal, captures absolute unfamiliarity: higher surprisal indicates greater deviation from dominant patterns and provides more informative learning signals.
 A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
@@
 Student- Agnostic Teacher Params 0.04 0.34 0.2 0.02 0.26 0.01
+Teacher Performance 0.49 0.34 0.13 0.23 0.03 0.23
 Avg-Token Length 0.49 0.68 0.45 0.57 0.47 0.53
-Teacher Performance 0.49 0.34 0.13 0.23 0.03 0.23
 Verified Accuracy 0.54 0.43 0.25 0.35 0.10 0.33
@@
 Our metric, Rank-Surprisal Ratio, achieves the highest correlation across all students.
-For each of the five student models, we compute average trajectory-level suitability (or quality) scores for reasoning datasets generated by eleven teacher models (§ 2.1) using different metrics.
-We then measure the correlation between these dataset-level scores and the student’s average post-training reasoning performance on the corresponding teacher data.
-We primarily report Spearman’s correlation coefficient, as Pearson correlation exhibits similar trends.
-We use a clipping threshold $r_{max}=100$ for our RSR metric in all subsequent experiments.
+For each of the five student models and each metric, we aggregate trajectory-level suitability (or quality) scores to obtain dataset-level scores for reasoning datasets generated by eleven teacher models (§ 2.1).
+We then measure the correlation between these dataset-level scores and the student’s reasoning performance after training on the corresponding teacher-generated datasets.
+We primarily report Spearman’s correlation coefficient, while Pearson correlation exhibits similar trends (§ C.2).
+For dataset-level RSR, we adopt a weighted averaging scheme (§ A.8) similar to Eq.
+6, which yields slightly higher correlation than a simple average of trajectory-level RSR.
+We use a clipping threshold of $r_{max}=100$ for RSR in all subsequent experiments.
 Additional analysis details are provided in § A.7, and complete metric scores are reported in § C.2.
@@
 We compare RSR against a diverse set of metrics for evaluating reasoning trajectories.
-These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal; [18]) and rank-based metrics.
+These include previously discussed teacher-side indicators (e.g., teacher model performance), basic statistics such as token length, as well as probability-based metrics (e.g., average surprisal and local surprisal [19]) and rank-based metrics.
 We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [45], LLM-judged quality scores, and answer accuracy on verifiable questions.
-In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE; [29]) and influence scores [16].
+In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE [30]) and influence scores [17].
 Results
@@
 […]
+[…]
    Rank clipping: $r_{max}=50$ $0.696$ 0.160-0.160
    Rank clipping: $r_{max}=500$ $0.822$ 0.034-0.034
-[…]
    Reduced sample size (200) $0.864$ $0.007$
 Table 5: Ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation.
-The derivation of Rank-Surprisal Ratio involves several design components as well as a hyperparameter $r_{max}$.
-We conduct an ablation study to examine how these choices affect correlation performance.
+The derivation of Rank-Surprisal Ratio involves several design components, as well as a hyperparameter $r_{max}$.
+We conduct an ablation study to examine how these choices affect the correlation strength of dataset-level RSR.
 As shown in Table 5, removing either rank clipping or the surprisal-weighted averaging substantially degrades the correlation, validating the necessity of both components in our metric.
@@
 The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$.
-More ablation results are provided in § C.3.
+Additional ablation results are in § C.3.
 5 Practical Applications
@@
 We then fine-tune student models on the constructed datasets and evaluate their post-training reasoning performance on standard math benchmarks, consistent with previous experiments.
-We compare RSR against a random selection baseline and multiple previously discussed metrics.
-For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the score.
+We compare RSR against a random selection baseline and several previously discussed metrics.
+For metric-based methods, candidate trajectories are scored and selected by either maximizing or minimizing the corresponding trajectory-level score.
 Results
@@
 Experimental Setting
-The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student.
-We consider a low-resource setting in which generating full training data for every teacher model is infeasible.
-Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the average trajectory-level score.
+The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation.
+We consider a realistic low-resource setting where generating full training data for every candidate teacher model is either costly or infeasible.
+Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the resulting dataset-level average score.
 We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@
 A.1 Determining Training Problem Set
-All teacher trajectory datasets are constructed using a fixed problem set to avoid confounding effects from variations in problem composition.
+All teacher trajectory datasets are constructed using a fixed problem set to avoid confounding effects from variations in problem composition, enabling a more controlled study.
 We focus on mathematical reasoning and curate 5,000 math problems from the widely used NuminaMath dataset [22].
-Following prior work [20], we apply preprocessing steps such as difficulty filtering to ensure the problem quality.
+Following prior work (Sky-T1) [21], we apply preprocessing steps such as difficulty filtering to ensure the problem quality.
 Specifically, our training dataset consists of problems drawn from the MATH, AIME/AMC, and Olympiads.
-SkyThought provides a difficulty-labeled version of NuminaMath in which each problem is assigned an integer difficulty score from 0 to 9.
+Sky-T1 provides a difficulty-labeled version of NuminaMath in which each problem is assigned an integer difficulty score from 0 to 9.
 Using this scale, we randomly sampled 1,667 problems from the MATH subset with difficulty above three, 1,667 problems from the Olympiads subset with difficulty above eight, and 1,666 problems from the AIME/AMC subset, yielding a balanced training set of 5,000 problems.
-We choose 5,000 training problems for three main reasons.
-First, prior studies have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems.
-Second, high-quality reasoning problems are relatively scarce, making further scaling less effective in practice.
-Third, our cross-study incurs quadratic computational cost across teacher–student pairs.
+We fix the training set size at 5,000 problems for three reasons.
+First, prior studies [45, 29] have shown that strong reasoning capabilities can be learned from training sets of around 1,000 problems.
+Second, high-quality reasoning problems are relatively scarce, making further scaling often unrealistic in practice.
+Third, our distillation study incurs quadratic computational costs across teacher–student pairs.
 Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
 A.2 Teacher Trajectory Generation
+For each teacher model, we generate reasoning trajectories for 5,000 problems over three independent runs using vLLM [20] under a maximum generation budget of 31,000 tokens.
+Each independent run (rollout) produces a dataset of 5,000 problem-trajectory pairs, yielding 11 × 3 = 33 datasets in total.
 […]
-If a sampled trajectory exceeds the token budget, we resample up to 10 times; if all attempts still exceed the budget, we truncate the final trajectory.
-Across all teacher models, truncation rates are below 1%, which helps preserve the quality of training trajectories even for teachers that tend to produce long outputs.
 […]
+We release all 33 generated trajectory datasets on Hugging Face. [^3] https://huggingface.co/datasets/Umean/RSR_data
 A.3 Teacher Models and Student Models
@@
 Table 8: List of teacher models used in our experiments.
-As discussed earlier, our experiments adopt a more diverse set of teacher models than prior work. Specifically, we consider the teacher variants shown in Table 8.
+As discussed earlier, our experiments adopt a more diverse set of teacher models than prior work.
+Specifically, we consider the teacher variants shown in Table 8.
+This set includes several recently emerged reasoning models, extending beyond the DeepSeek-R1 and QwQ teachers commonly used in prior studies.
+These teachers produce trajectories that vary in length, informational content, and style.
+For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed.
+Concrete trajectory examples from different teachers can be inspected in our Hugging Face datasets referenced above.
 For student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B.
@@
 Table 9: Training hyperparameters for different student models.
-We use LLaMA-Factory [50] framework to conduct SFT with reasoning trajectory datasets.
+We perform SFT on reasoning trajectory datasets using the LLaMA-Factory [52] framework and FlashAttention-2 [8].
 For different student models, we conduct grid search for best set of hyperparameters.
@@
 During SFT, all student models use a maximum sequence length of 32,768.
-Most experiments are conducted on NVIDIA H200 GPUs. A single SFT experiment using the 14B model takes approximately 6 hours on 8 H200 GPUs for training.
+Most experiments are conducted on NVIDIA H200 GPUs. A single SFT experiment using the 7B model takes around 7 hours on 8 H200 GPUs for training.
 For trajectory selection experiments in § 5.1, we fine-tune each selected dataset with three different random seeds and report performance averaged over these runs.
@@
 We use the Spearman correlation coefficient [47] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance.
-We also report Pearson correlation results [7] in § C.2.
+We also report Pearson correlation [7] results in § C.2.
 We report the absolute values of Spearman correlation coefficients in Table 4.
@@
 We observe that our datset-level metric varies marginally across different datasets generated by the same teacher.
+A.8 Definition of Dataset-level RSR
+For dataset-level RSR, we adopt a surprisal-weighted averaging scheme analogous to the trajectory-level weighted average (§ 3.4).
+This design aims to mitigate numerical instability caused by trajectories with disproportionately small average surprisal and to emphasize trajectories with larger average surprisal during dataset-level aggregation.
+We now formally define the dataset-level RSR used in the correlation analysis (§ 4.1, 4.2).
+Let $\mathbf{X}=\{\mathbf{x}_{j}\}_{j=1}^{|\mathbf{X}|}$ denote a dataset of trajectories, where each trajectory $\mathbf{x}_{j}$ contains response tokens $\{t_{j,k}\}_{k=1}^{|\mathbf{x}_{j}|}$.
+For brevity, we define the trajectory-level average clipped rank and average surprisal as
+$\overline{r}(\mathbf{x}_{j})\;=\;\frac{1}{|\mathbf{x}_{j}|}\sum_{k=1}^{|\mathbf{x}_{j}|}\min\!\bigl(\mathrm{Rank}(t_{j,k}),r_{max}\bigr),\qquad\overline{s}(\mathbf{x}_{j})\;=\;\frac{1}{|\mathbf{x}_{j}|}\sum_{k=1}^{|\mathbf{x}_{j}|}\mathrm{Surprisal}(t_{j,k}).$ (8)
+Accordingly, the trajectory-level RSR can be equivalently written as $\mathrm{RSR}(\mathbf{x}_{j})=\frac{\overline{r}(\mathbf{x}_{j})}{\overline{s}(\mathbf{x}_{j})}$.
+To obtain the dataset-level RSR, we apply a weighted averaging scheme analogous to Eq. 6, using the trajectory-level average surprisal $\overline{s}(\mathbf{x}_{j})$ as the weight:
+$\mathrm{RSR}_{\text{dataset}}(\mathbf{X})$ $=\;\frac{\sum_{j=1}^{|\mathbf{X}|}\overline{s}(\mathbf{x}_{j})\,\mathrm{RSR}(\mathbf{x}_{j})}{\sum_{j=1}^{|\mathbf{X}|}\overline{s}(\mathbf{x}_{j})}$ (9)
+$=\;\frac{\sum_{j}\overline{s}(\mathbf{x}_{j})\,\frac{\overline{r}(\mathbf{x}_{j})}{\overline{s}(\mathbf{x}_{j})}}{\sum_{j}\overline{s}(\mathbf{x}_{j})}$
+$=\;\frac{\sum_{j}\overline{r}(\mathbf{x}_{j})}{\sum_{j}\overline{s}(\mathbf{x}_{j})}.$
+The resulting dataset-level metric admits a concise form, which can be interpreted as the ratio of the summed trajectory-level average clipped ranks to the summed trajectory-level average surprisals.
+We compare the correlation strength of the simple dataset-level average of trajectory-level RSR, i.e., $\frac{1}{|\mathbf{X}|}\sum_{j=1}^{|\mathbf{X}|}\mathrm{RSR}(\mathbf{x}_{j})$, with that of the dataset-level RSR computed via surprisal-weighted averaging (Eq.
+9) in the additional ablation study (§ C.3).
 Appendix B Details of Compared Metrics
@@
 We adopt the Influence Score method based on the second-order approximation of influence functions [17].
-The core principle is to treat the infinitesimal up-weighting of a training sample as a local perturbation to the optimization objective, thereby estimating the induced marginal change in the loss of the evaluation set.
-In this framework, a higher influence score signifies that the gradient direction of the training sample aligns with minimizing the evaluation loss in the neighborhood of the converged parameters.
+The core idea is to model an infinitesimal up-weighting of a training sample as a local perturbation to the optimization objective, thereby estimating the resulting marginal change in the evaluation-set loss.
+Within this framework, a higher influence score indicates that the gradient direction induced by the training sample is more aligned with minimizing the evaluation loss in the vicinity of the converged parameters.
 We implement this baseline using the Kronfluence framework.
@@
 To optimize for memory and computational efficiency, we apply a low-rank approximation to the query gradients (rank =4=4) and utilize bfloat16 precision for the Inverse Hessian-Vector Product calculations.
-Finally, we average these pairwise scores across the evaluation set to obtain the sample-level influence score $s(d)$, which serves as our baseline for ranking data suitability.
+Finally, we average these pairwise scores across the evaluation set to obtain the sample-level (trajectory-level) influence score $s(d)$, which is used as the baseline for evaluating data suitability.
 B.4 G-Norm
@@
 To address the computational constraints associated with high-dimensional parameter spaces, the method utilizes random projection for dimensionality reduction.
-For each reasoning trajectory $(x,y)$, we compute the gradient of the loss derived from the student model and apply length normalization to eliminate bias towards longer sequences.
+For each reasoning trajectory $\mathbf{x}$, we compute the gradient of the loss derived from the student model and apply length normalization to eliminate bias towards longer sequences.
 These gradients are then projected onto a lower-dimensional subspace via a fixed random matrix, followed by the computation of their $L_{2}$ norms.
@@
 B.6 Others
-Here we give formal definitions for Avg-Surprisal and Avg-Surplocal. Given a trajectory $x=(t_{1},\ldots,t_{T})$,
+Here we give formal definitions for Avg-Surprisal and Avg-Surplocal3.2). Given a trajectory $x=(t_{1},\ldots,t_{T})$,
 $\mathrm{Avg\_Surprisal}(\mathbf{x})=\frac{1}{T}\sum_{k=1}^{T}-\log p_{\theta}(t_{k}\mid\mathbf{c}_{k})$ (10)
@@
 C.1 Computational Cost for Rank-Surprisal Ratio
-The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model.
-During the forward pass, we collect each token’s surprisal and rank from the model logits.
-Computing Rank-Surprisal Ratio over the 5,000-trajectory dataset typically takes about one hour on a single H200 GPU, which is significantly cheaper than SFT.
+The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model, during which we collect each token’s surprisal and clipped rank from the model logits.
+For a single trajectory, computing token-level surprisals and ranks has a worst-case time complexity of $\mathcal{O}(TV)$, where $T$ denotes the number of response tokens and $V$ is the vocabulary size.
+This computation does not introduce additional complexity beyond that inherent in the dimensionality of the model logits.
+Moreover, since RSR only depends on clipped ranks, the rank computation can be efficiently implemented via top-$k$ selection rather than full-vocabulary comparisons, reducing practical runtime and GPU memory usage.
+In practice, computing RSR over the 5,000-trajectory dataset with a context length of 32,768 using a 7B model typically takes under one hour on a single H200 GPU with FlashAttention-2, which is significantly cheaper than SFT.
 C.2 Additional Results for Correlation Analysis
-Metrics Student Models
+Metric Correlation Measure Student Models
 Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
@@
 Table 10: Comparison of Spearman and Pearson correlation results on Rank-Surprisal Ratio (absolute values).
-Table 10 compares Spearman and Pearson correlations. The results show that Rank-Surprisal Ratio also exhibits strong Pearson correlation with post-training performance.
+Table 10 reports both Spearman and Pearson correlations for Rank-Surprisal Ratio.
+The results show that RSR also exhibits strong Pearson correlation with post-training performance.
+The rationale for primarily using Spearman correlation rather than Pearson correlation is discussed in § A.7.
 Table 22, 23, 24, 25, and 26 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
@@
    Fixed student $0.785$ 0.071-0.071
+   Simple dataset-level average $0.838$ 0.018-0.018
    Use Rank1.05 $0.845$ 0.011-0.011
@@
    Use Surprisal0.95 $0.873$ $0.017$
-Avg-Rank (clip at $r_{max}=100$) $0.552$ 0.304-0.304
+Avg-Rank (clipped) $0.552$ 0.304-0.304
 Rank Minus Surprisal $0.585$ 0.271-0.271
@@
 The “Fixed student” variant computes RSR using a fixed model (Qwen-3-14B) instead of the target student, and the resulting drop in correlation highlights the importance of student-specific estimation.
-“Avg-Rank (clip at $r_{max}=100$)” applies rank clipping to Avg-Rank, but it still fails to achieve better correlation than rank alone.
-“Rank Minus Surprisal” is computed by subtracting the surprisal from the (clipped) rank.
-Although this metric also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance.
-Rank–Entropy Ratio is computed as the ratio between the (clipped) average rank and the average token entropy, where entropy is loosely related to surprisal.
-While it achieves good correlation, its correlation is still lower than that of RSR.
-We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to power 1 for both rank and surprisal), for example "use Rank1.05".
+The "Simple dataset-level average" variant computes the dataset-level RSR by simply averaging trajectory-level scores, as noted in § A.8.
+The slight degradation in correlation indicates that a simple average of trajectory-level RSR remains a robust estimator of dataset-level suitability, while the surprisal-weighted averaging scheme improves the reliability of the aggregated metric at both trajectory-level and dataset-level.
+The “Avg-Rank (clipped)” variant applies rank clipping with $r_{max}=100$ to Avg-Rank; however, it still exhibits a notable performance gap compared with RSR.
+This suggests that rank information alone is insufficient to yield strong correlation, underscoring the importance of jointly leveraging rank and surprisal.
+The “Rank Minus Surprisal” metric is computed by subtracting the surprisal from the (clipped) rank.
+Although it also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance.
+"Rank–Entropy Ratio" is defined as the ratio between the average (clipped) token rank and the average token entropy, where entropy is loosely related to surprisal.
+While this metric achieves relatively strong correlation, its performance remains below that of RSR.
+We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to a power of 1 for both rank and surprisal), for example "use Rank1.05".
 The results indicate that our metric is robust to these variations and may yield higher correlation with hyperparameter tuning.
@@
 Selected Datasets Data Composition over Teacher Models Metrics
-Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR
-RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57
-RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69
-RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67
-RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56
-RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73
-Table 14: Data composition and metrics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) across different student models.
+R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR Length
+RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57 9363
+RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69 9939
+RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67 9845
+RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56 9419
+RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73 10169
+Table 14: Data composition and metric statistics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) for different student models.
 Model names are abbreviated as Q for Qwen and L for LLaMA.
@@
 For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed.
-The RSR values of the selected datasets are also reported in Table 14.
+The average RSR values of the selected datasets are also reported in Table 14.
 These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § C.2 for complete metrics), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
 C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
+To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers.
+Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B.
+This teacher set is formed by combining the teachers used in Table 2 and 7.
 Selection Methods Qwen-3-14B Qwen-2.5-7B
@@
 Rank-Surprisal Ratiomin 76.8 50.0
-Table 15: Comparison of different trajectory selection methods under a reduced-teacher setting.
-To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers.
-Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B.
-This teacher set is formed by combining the teachers used in Table 2 and 7.
+Table 15: Comparison of different trajectory selection methods under a reduced-teacher setting. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
 The results are shown in Table 15.
@@
 Avg-Rank 49.412 55.609 430.379 65.293 55.249 58.448 41.968 365.886 303.840 45.098 49.145
+Avg-Rank (clipped) 1.930 1.811 4.097 1.422 1.682 1.728 1.355 4.652 3.589 1.457 1.695
 GRACE 0.028 0.028 0.168 0.025 0.031 0.030 0.025 0.154 0.113 0.023 0.026
@@
 Avg-Rank 10.328 11.101 72.356 11.762 11.028 11.206 8.463 66.664 49.643 8.721 9.757
+Avg-Rank (clipped) 2.831 2.821 5.633 2.183 2.687 2.665 2.016 6.178 4.638 2.174 2.552
 GRACE 0.162 0.177 1.005 0.185 0.183 0.181 0.111 0.853 0.639 0.143 0.159
@@
 Avg-Rank 6.192 6.438 36.628 6.413 6.312 6.411 4.724 35.233 25.818 5.000 5.683
+Avg-Rank (clipped) 2.477 2.416 4.557 1.842 2.280 2.264 1.710 5.189 3.921 1.869 2.198
 GRACE 0.168 0.199 1.760 0.126 0.160 0.142 0.119 1.491 0.908 0.120 0.139
@@
 Avg-Rank 9.172 10.267 82.658 10.747 9.773 10.346 7.503 72.151 54.562 7.710 9.011
+Avg-Rank (clipped) 2.132 2.023 4.581 1.526 1.835 1.896 1.422 5.130 3.955 1.535 1.802
 GRACE 0.148 0.151 0.524 0.148 0.165 0.145 0.139 0.436 0.425 0.124 0.116
@@
 Avg-Rank 18.247 19.034 145.459 24.048 21.422 19.906 16.483 119.646 90.940 16.967 17.978
+Avg-Rank (clipped) 2.794 2.747 5.307 2.040 2.550 2.551 1.855 5.870 4.381 2.049 2.462
 GRACE 0.096 0.092 0.602 0.093 0.100 0.147 0.077 0.991 0.426 0.075 0.085

### v3

@@
 A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
+§ C.5 presents illustrative examples of RSR measurements.
 4 Correlation Analysis
@@
 Rank-Surprisal Ratiomin 67.5 59.2 93.1 94.6 78.6 28.5 53.2 61.4 34.8
-Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods.
+Table 6: Trajectory selection results showing post-training reasoning performance of student models trained on datasets selected by different methods (5k samples each).
 maxmax and minmin indicate maximizing or minimizing the corresponding metric.
 Model names are abbreviated as Q for Qwen and L for LLaMA.
-Additional results, including GPQA evaluation, are provided in § C.4.
+Additional results, including the GPQA evaluation and the No-Selection baseline, are provided in § C.4.
 Experimental Setting
 The trajectory selection task aims to identify the most effective reasoning trajectory from a set of candidates for a given problem or prompt.
-We adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion.
+This aligns with the widespread need for data-efficient training in real-world scenarios, particularly when resources are limited.
+Specifically, we adopt a 33-to-1 setting, where each candidate pool contains 33 trajectories generated by 11 teacher models (3 per teacher; see § 2.1), and one trajectory is selected according to the selection criterion.
 This procedure is repeated for all 5,000 training problems and all student models, yielding student-specific 5,000-trajectory teacher datasets for each selection method.
@@
 These results underscore the practical value of RSR for selecting effective trajectories prior to training.
-Additional trajectory selection results, including complete scores, further analysis, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
+Additional trajectory selection results, including complete scores, further analysis, the No-Selection baseline, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
 5.2 Teacher Selection
@@
 Experimental Setting
-The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation.
-We consider a realistic low-resource setting where generating full training data for every candidate teacher model is either costly or infeasible.
-Instead, we sample a small set of 200 trajectories from each candidate teacher, score them using different metrics, and select the teacher model based on the resulting dataset-level average score.
+The teacher selection task aims to identify the most suitable teacher model for generating reasoning trajectories to train a given student model prior to distillation, reflecting a recurring challenge in real-world applications.
+To better capture practical constraints, we consider a low-resource setting in which generating full training data for every candidate teacher model is either costly or infeasible.
+Instead, we sample a small set of 200 trajectories from each candidate teacher, evaluate them using different metrics, and select the teacher model based on the resulting dataset-level average score.
 We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@
 Overall, our results highlight informative alignment as a promising direction for reasoning distillation.
+8 Broader Applicability of RSR
+Beyond CoT Data
+Although our work focuses solely on reasoning tasks, RSR is not specifically designed for reasoning trajectories (CoTs) and can be applied to general text data.
+Exploring its behavior and effectiveness beyond reasoning settings remains an interesting direction for future work.
+Subset Selection
+Beyond trajectory and teacher selection, RSR may also be applicable to subset selection.
+Subset selection aims to identify a high-quality subset from a dataset composed of different reasoning problems and their corresponding trajectories.
+Unlike trajectory selection, which compares multiple trajectories for the same problem, subset selection requires cross-problem comparison of RSR values to filter samples across heterogeneous reasoning prompts.
+This scenario is particularly relevant when only a single trajectory is available per problem, yet improved data efficiency is still desired.
+Although subset selection introduces additional confounding factors due to variation across problems—making it harder to isolate trajectory–student suitability effects—RSR may still help identify relatively more effective samples across different reasoning problems.
+Preliminary experiments on internal data provide suggestive evidence for this possibility.
 Limitations
@@
 […]
-We release all 33 generated trajectory datasets on Hugging Face. [^3] https://huggingface.co/datasets/Umean/RSR_data
+We release all 33 generated trajectory datasets along with RSR-selected subsets (§ 5.1) on Hugging Face. [^3] https://huggingface.co/datasets/Umean/RSR_data
 A.3 Teacher Models and Student Models
@@
 For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed.
-Concrete trajectory examples from different teachers can be inspected in our Hugging Face datasets referenced above.
-For student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B.
+Specific trajectory examples from different teachers are available in our released datasets on Hugging Face and in our case study (§ C.5).
+For the student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B.
+We use base models instead of chat models as students, following the training practice of recent reasoning models.
+This choice ensures a clean starting point with greater potential [41], allows us to observe more substantial training effects, and avoids stylistic mismatch and overlapping supervision between instruction-tuning data and reasoning trajectories.
 A.4 Benchmark Evaluation
@@
 Most experiments are conducted on NVIDIA H200 GPUs. A single SFT experiment using the 7B model takes around 7 hours on 8 H200 GPUs for training.
-For trajectory selection experiments in § 5.1, we fine-tune each selected dataset with three different random seeds and report performance averaged over these runs.
-For teacher–student distillation experiments in § 2.1, however, each teacher already yields three independent trajectory datasets, so we do not further vary random seeds for fine-tuning.
+For trajectory selection experiments in § 5.1, we fine-tune each selected dataset with three random seeds and report performance averaged over these runs.
+For distillation experiments in § 2.1, however, each teacher already yields three independent trajectory datasets, so we do not further vary random seeds for fine-tuning.
 A.6 Details of Simulation Study
@@
 The rationale for primarily using Spearman correlation rather than Pearson correlation is discussed in § A.7.
-Table 22, 23, 24, 25, and 26 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
+Table 23, 24, 25, 26, and 27 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
 C.3 Additional Ablation Study
@@
 C.4 Additional Results for Trajectory Selection
-C.4.1 Ablation Study on Trajectory Selection
+C.4.1 RSR-based Trajectory Selection Variants and No-Selection Baseline
 Variant Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
 Math Avg. Math Avg. Math Avg. Math Avg. Math Avg.
-Rank-Surprisal Ratiomin 78.6 28.5 53.2 61.4 34.8 51.3
+Rank-Surprisal Ratiomin (5k) 78.6 28.5 53.2 61.4 34.8 51.3
    With Correctness Filtering 77.5 27.9 52.3 60.7 34.9 50.7
    Fewer Candidates per Teacher 77.6 27.8 52.6 60.9 33.8 50.5
-Table 12: Ablation study of trajectory selection with RSRmin-based variant methods across student models. "Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
-Table 12 presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
+No Selection (55k combined) 72.7 41.5 54.4 58.9 35.3 52.5
+Table 12: Performance of RSR-based trajectory selection variants and a No-Selection baseline trained on a 55k combined trajectory dataset.
+"Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
+Although trajectory selection enables data-efficient training and offers substantial practical value under resource-constrained settings, we are also interested in how it compares with training on all available data without any selection when resources are sufficient.
+To this end, we construct a “No Selection” baseline.
+Unlike RSR, which selects 5k trajectories from different teachers, this baseline directly merges 11 trajectory datasets—one per teacher—into a unified training set comprising 55k samples.
+We do not use the full 165k dataset (i.e., all three trajectory datasets per teacher) for two reasons.
+First, the computational cost would be prohibitively high.
+Second, training on multiple highly similar trajectory datasets from the same teacher may reduce data diversity and increase the risk of overfitting.
+We consider the 55k setting to be sufficiently representative and more consistent with practical use cases.
+From Table 12, we observe that the 5k trajectories selected by RSRachieve post-training performance that is comparable to—or even exceeds—the 55k-sample “No Selection” baseline on most student models.
+An exception is LLaMA-3.1-8B, which may require larger-scale data to develop reasoning ability due to its relatively limited prior exposure to reasoning tasks.
+In contrast, for stronger students (Qwen-3-14B and Qwen-3-4B), the selected 5k trajectories significantly outperform the unfiltered 55k dataset.
+This suggests that a small amount of high-quality data is sufficient to effectively activate the reasoning capabilities of student models.
+Overall, these findings demonstrate that trajectory selection not only improves data efficiency but can also deliver superior performance with substantially fewer resources, further highlighting the practical value of RSR.
+Table 12 also presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
 For "With Correctness Filtering", we consider a combined setting that uses both RSR and verified correctness for trajectory selection.
@@
 This suggests that RSR can identify suitable trajectories that consistently improve student models’ general reasoning capabilities, even when training solely on mathematical problems.
-The complete trajectory selection results underlying Table 6 are presented in Table 17, 18, 19, 20, and 21.
+The complete trajectory selection results underlying Table 6 are presented in Table 18, 19, 20, 21, and 22.
 C.4.3 Analysis of Datasets Selected by RSR
@@
 For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed.
+We release the RSR-selected datasets for five student models on Hugging Face.
 The average RSR values of the selected datasets are also reported in Table 14.
@@
 Table 16: Data composition of trajectory datasets selected by RSR under a reduced-teacher setting (see the setting in § C.4.4 and results in Table 15).
+C.5 What Does RSR Prefer? A Token-Level Case Study
+To better understand how Rank-Surprisal Ratio operates in practice and what kinds of trajectories it prefers, we conduct a token-level case study comparing trajectories generated by different teacher models.
+[…]
+As shown, tokens in trajectories from GPT-OSS-20B often exhibit excessively high rank values, suggesting that these trajectories are relatively unfamiliar to the student model and misaligned with the student’s current behavior.
+This may stem from the fast-paced reasoning flow and uncommon phrasing in GPT-OSS-20B trajectories, which make them harder for the student to follow.
+Meanwhile, tokens from Nemotron-Super frequently have very low surprisals, indicating that these trajectories closely resemble the student’s own generations and therefore provide limited informativeness.
+This may be attributed to the relatively conventional reasoning flow and limited analytical depth of Nemotron-Super’s trajectories.
+As a result, tokens from both models tend to receive high token-level RSR values, implying less effective supervision for student training.
+By comparison, trajectories from QwQ-32B exhibit a moderate level of comprehensibility and depth.
+Their tokens show relatively lower rank values alongside higher surprisals, striking a better balance between alignment and informativeness.
+Consequently, they yield lower token-level RSR values and are preferred under the RSR criterion.
+These observations are consistent with the distillation results (Table 1) and the subsequent analysis presented in the main body of the paper.
+We also observe that token-level RSR may become numerically unstable when extremely low surprisal yields inflated RSR values (e.g., 100+).
+This motivates our surprisal-weighted averaging for trajectory-level RSR (§ 3.4), which stabilizes the metric and improves its effectiveness.
+Additional trajectory examples from 11 teachers are available in our released datasets on Hugging Face.
+[…]
+Table 17: A case study with visualization of token-level rank, surprisal, and RSR values (with rank clipping) across trajectories generated by GPT-OSS-20B, Nemotron-Super-49B-v1.5, and QwQ-32B, as measured by the Qwen-2.5-7B student.
+Each token is color-coded according to its RSR value, where darker red indicates lower values (stronger preference under RSR).
+GPT-OSS-20B exhibits uncommon phrasing and high rank values; Nemotron-Super exhibits conventional reasoning flow and low surprisal; QwQ-32B strikes a more favorable balance between informativeness and alignment, yielding lower token-level RSR values.
+Trajectory-level RSR is defined as the surprisal-weighted average of token-level RSR values with rank clipping (§ 3.4).
 Appendix D Others
@@
 Table 27: Full metric assessment results on Qwen-2.5-3B. Model names are abbreviated with Q for Qwen.
-Prompt
+Evaluation prompt for LLM-judged quality assessment
 Table 28: Evaluation prompt for LLM-judged quality assessment.

### v4

@@
 Motivated by our preliminary analysis, we argue that the dilemma between providing new signals and aligning with student’s existing behavior can be resolved by trajectories exhibiting both absolute unfamiliarity and relative familiarity.
-Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while still containing behavioral patterns within the student’s prior experience, such that their tokens rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
-Figure 1: Illustration of the intuition behind RSR.
+Concretely, effective trajectories should deviate from the student’s own generations, receiving low absolute probability under the student model, while remaining compatible with its overall generation patterns, such that their tokens still rank relatively high in the model’s prediction distribution over the vocabulary (Figure 1).
+Figure 1: Illustration of the intuition behind Rank-Surprisal Ratio.
 Suitable reasoning trajectories should balance informativeness and alignment, having low absolute probability while their tokens remain relatively high-ranked under the student model.
@@
 Our main contributions are three-fold:
-  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data-student suitability.
-  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory for a given student model by jointly capturing alignment and informativeness, achieving a strong correlation with post-training performance.
-  • We demonstrate the practical utility of Rank-Surprisal Ratio in two data engineering scenarios, trajectory selection and teacher selection, where it serves as an effective criterion and outperforms existing methods.
+  • We present a systematic distillation study across a wide range of teacher and student models, showing that the effectiveness of reasoning trajectories differs across students and highlighting the importance of data-student suitability (§ 2).
+  • We propose Rank-Surprisal Ratio, a simple metric that quantifies the suitability of a reasoning trajectory for a given student model by jointly capturing alignment and informativeness (§ 3), achieving a strong correlation with post-training performance (§ 4).
+  • We demonstrate the practical utility of RSR in two data engineering scenarios, trajectory selection and teacher selection, where it serves as an effective criterion and outperforms existing methods (§ 5).
+We further discuss the broader applicability of RSR to non-CoT data and subset selection (§ 7).
 2 The Need for Student-Specific Data
-To understand which types of reasoning trajectories most effectively improve student models after SFT, we conduct a comprehensive large-scale study involving five widely adopted student models and eleven diverse reasoning-oriented teacher models, yielding 55 teacher–student pairings.
+To understand which types of reasoning trajectories most effectively improve student models after SFT, we conduct a comprehensive large-scale study involving five widely adopted student models and eleven diverse reasoning-oriented teacher models, yielding 55 teacher-student pairings.
 We perform SFT experiments for each of these pairs.
@@
 Darker and lighter shading indicate the best and second-best results, respectively.
-Student performance varies significantly across teacher–student pairs, highlighting the importance of data–student suitability.
+Student performance varies significantly across teacher-student pairs, highlighting the importance of data-student suitability.
 2.1 Experimental Settings
 Our teacher-student pairing study involves two major steps: (1) For each teacher model, we prompt it to generate a long CoT response for each math problem in our 5000-problem set (see § A.1), forming a trajectory dataset specific to that teacher.
-(2) For each teacher–student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance.
+(2) For each teacher-student pair, we fine-tune the student model on the corresponding teacher dataset and evaluate its reasoning performance.
 All students are pre-trained base models.
-To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher–student pair.
+To reduce variance induced by stochastic trajectory sampling, we perform three independent generation runs for each teacher and conduct SFT separately on each resulting dataset for every teacher-student pair.
 Reported results are averaged over these three runs.
@@
 2.2 Results
-Table 1 presents the results of our teacher–student pairing distillation study, revealing that:
+Table 1 presents the results of our teacher-student pairing distillation study, revealing that:
 Stronger teachers do not necessarily produce better students.
@@
 Similarly, teachers with strong reasoning performance do not consistently yield the best outcomes for all student models.
-Data–student suitability is critical for eliciting reasoning improvements.
+Data-student suitability is critical for eliciting reasoning improvements.
 The effectiveness of teacher trajectories is highly student-specific and depends critically on their suitability for the student model.
@@
 Moreover, teachers from distant model families (GPT-OSS) often lead to inferior results, suggesting that unfamiliar reasoning patterns are harder for students to absorb.
-Overall, we find no simple teacher–student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
+Overall, we find no simple teacher-student pairing rule based on surface attributes such as parameter scale or model family, indicating that reasoning data suitability is a nuanced property requiring deeper investigation.
 3 Measuring Data-Student Suitability
-In this section, we explore metrics for measuring data–student suitability, with the goal of jointly capturing informativeness and alignment.
+In this section, we explore metrics for measuring data-student suitability, with the goal of jointly capturing informativeness and alignment.
 We begin by introducing two fundamental token-level measures: surprisal and rank (§ 3.1), and analyzing the limitations of existing probability-based metrics (§ 3.2).
@@
 […]
-Table 2: Comparison of the student’s post-training performance and data–student suitability metrics across trajectories from different teacher models, evaluated on Qwen-2.5-7B.
+Table 2: Comparison of the student’s post-training performance and data-student suitability metrics across trajectories from different teacher models, evaluated on Qwen-2.5-7B.
 Darker shading indicates higher performance or better suitability.
 Metrics whose trends align with performance (e.g., RSR) provide more reliable suitability estimates.
-Complete metric scores are provided in § C.2.
+Complete metric scores for all teacher and student models are provided in § F.
 3.1 Surprisal and Rank
@@
 Conversely, alignment does not entail exact agreement with the student’s outputs, but instead requires that the corresponding tokens have relatively higher likelihoods than other candidates in the vocabulary.
-Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while still exhibiting behavioral patterns within the student’s prior experience.
+Building on this insight, we propose that effective reasoning trajectories should deviate from the student’s own generations while remaining compatible with the student’s overall generation patterns learned from prior experience.
 Therefore, tokens in such trajectories are assigned low absolute probability (i.e., high surprisal) by the student model while still ranking relatively high (i.e., having low rank values) in its prediction distribution.
@@
 Specifically, we model the student’s prediction distribution as a bimodal distribution over the vocabulary $\mathcal{V}$.
-The first mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant patterns, which arise from student’s massive general training data.
-The second mode $Z_{B}$, represents tokens whose patterns resemble strong reasoning trajectories, reflecting the student’s prior exposure to relevant data.
+The first (major) mode, denoted as $Z_{A}$, represents tokens that follow the student’s dominant generation patterns, which arise from its general training data.
+The second (minor) mode, $Z_{B}$, occupies a small fraction of the probability mass.
+It represents tokens from specific generation patterns that deviate from the major mode yet remain familiar to the student model.
 We instantiate both $Z_{A}$ and $Z_{B}$ as Zipf distributions [56, 27] over $\mathcal{V}$.
@@
 where $\pi=\frac{M_{A}}{M_{A}+M_{B}}$ and $M_{A}>M_{B}$.
-Based on $Z$, we simulate four types of reasoning trajectories and examine their average token surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior.
+Based on $Z$, we simulate four types of reasoning trajectories and examine their average token surprisal and rank: (i) $X_{A}$, sampled from $Z_{A}$, representing trajectories that closely follow the student’s dominant generation patterns; (ii) $X_{B}$, sampled from $Z_{B}$, representing trajectories that deviate from the dominant mode while aligning with certain minor patterns within the student; (iii) $X_{C}$, sampled from a distribution distinct from both $Z_{A}$ and $Z_{B}$, representing misaligned trajectories; and (iv) $X_{D}$, sampled from $Z$, representing trajectories that reflect the student’s overall predictive behavior.
 More simulation details are provided in § A.6.
-Figure 2: Simulation of the student model’s token-level prediction distribution $Z$.
-Tokens with low probability but relatively high rank—characteristic of the $Z_{B}$ mode—yield smaller token-level rank–surprisal ratios.
+Figure 2: Simulation of the student model’s token-level bimodal prediction distribution $Z$.
+Tokens with low probability yet ranked near the top (typical of the minor mode $Z_{B}$) tend to yield smaller, and thus preferred, rank-surprisal ratios.
 […]
-$X_{A}$ (from $Z_{A}$) 0.41 1.38 2.49 1.69
-$X_{B}$ (from $Z_{B}$) 0.10 2.73 4.31 1.30
-$X_{C}$ (from $Z_{C}$) 0.08 4.73 11.57 2.23
-$X_{D}$ (from $Z$) 0.35 1.67 2.93 1.62
+$X_{A}$ (from $Z_{A}$, major) 0.41 1.38 2.49 1.69
+$X_{B}$ (from $Z_{B}$, minor) 0.10 2.73 4.31 1.30
+$X_{C}$ (from $Z_{C}$, misaligned) 0.08 4.73 11.57 2.23
+$X_{D}$ (from $Z$, mixed) 0.35 1.67 2.93 1.62
 […]
@@
 A natural solution is to exclude tokens with very low surprisal when computing the average.
-Let $\mathcal{T}_{H}(\mathbf{x})$ denote the set of response tokens whose surprisal lies in the top $H\%$ within a trajectory $\mathbf{x}$.
+Let $\mathcal{T}_{H}(\mathbf{x})$ denote the set of tokens whose surprisal lies in the top $H\%$ within a trajectory $\mathbf{x}$.
 We define a filtered average as
@@
 Empirically, we find that using the top 30% highest-surprisal tokens yields stronger correlation with post-training performance (Table 2).
-This suggests that tokens with higher surprisal have a greater impact on student learning and should therefore be emphasized when computing the average.
+This suggests that response tokens with higher surprisal have a greater impact on student learning and should therefore be emphasized when computing the average.
 Accordingly, instead of hard filtering, we adopt a surprisal-weighted average of the token-level ratios.
@@
 $\frac{\sum_{k}s_{k}\,\mathrm{RSR}_{\text{token}}(t_{k})}{\sum_{k}s_{k}}=\frac{\sum_{k}s_{k}\,\frac{r_{k}}{s_{k}}}{\sum_{k}s_{k}}\\ =\frac{\sum_{k}\mathrm{Rank}(t_{k})}{\sum_{k}\mathrm{Surprisal}(t_{k})}$ (6)
-The resulting metric yields a concise form that can be interpreted as the ratio of the average token rank to the average token surprisal over a trajectory.
+The resulting metric yields a concise form, interpretable as the ratio of the average token rank to the average token surprisal over a trajectory’s response tokens.
 In practice, we further observe that extremely unfamiliar tokens can attain very large rank values due to the large vocabulary size, which also leads to numerical instability.
-Since tokens with excessively large ranks are effectively indistinguishable from the student’s perspective, we clip rank values at a threshold $r_{max}$.
-Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR [^2] By default, RSR refers to the trajectory-level RSR in this paper, unless otherwise specified (e.g., in correlation analysis).), as
+Since such tokens are effectively indistinguishable to the student, we clip rank values at a threshold $r_{max}$ to improve stability without sacrificing meaningful information.
+Thus, we define our final trajectory-level metric, Rank-Surprisal Ratio (RSR) [^2] By default, RSR refers to the trajectory-level RSR in this paper, unless otherwise specified (e.g., in correlation analysis)., as
 $\mathrm{RSR}(\mathbf{x})\;=\;\frac{\sum_{k}\min\!\bigl(\mathrm{Rank}(t_{k}),r_{max}\bigr)}{\sum_{k}\mathrm{Surprisal}(t_{k})}$ (7)
@@
 A lower RSR therefore identifies trajectories that better balance alignment and informativeness, corresponding to effective reasoning supervision.
-§ C.5 presents illustrative examples of RSR measurements.
+§ D.2 presents illustrative examples of RSR measurements.
 4 Correlation Analysis
@@
 Rank-Surprisal Ratio 0.85 0.85 0.92 0.82 0.85 0.86
-Table 4: Spearman correlation between different metrics and post-training reasoning performance across student models, reporting absolute values.
-"Student-Agnostic" metrics are computed independently of the specific student model.
-Our metric, Rank-Surprisal Ratio, achieves the highest correlation across all students.
+Table 4: Spearman correlation between each data-student suitability metric and post-training reasoning performance (average accuracy on reasoning benchmarks), reported in absolute values for different student models.
+“Student-agnostic” metrics are computed independently of the specific student model.
 For each of the five student models and each metric, we aggregate trajectory-level suitability (or quality) scores to obtain dataset-level scores for reasoning datasets generated by eleven teacher models (§ 2.1).
 We then measure the correlation between these dataset-level scores and the student’s reasoning performance after training on the corresponding teacher-generated datasets.
-We primarily report Spearman’s correlation coefficient, while Pearson correlation exhibits similar trends (§ C.2).
+We primarily report Spearman’s correlation coefficient, while Pearson correlation exhibits similar trends (§ E.1).
 For dataset-level RSR, we adopt a weighted averaging scheme (§ A.8) similar to Eq.
@@
 We use a clipping threshold of $r_{max}=100$ for RSR in all subsequent experiments.
-Additional analysis details are provided in § A.7, and complete metric scores are reported in § C.2.
+Additional analysis details are provided in § A.7, and complete metric scores are reported in § F.
 Compared Metrics
@@
 We also consider commonly used trajectory quality measures, such as rule-based quality scores derived from word frequency [46], LLM-judged quality scores, and answer accuracy on verifiable questions.
-In addition, we include recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE [30]) and influence scores [17].
+In addition, we include other recent student-specific data suitability metrics, including gradient-based scores (G-Norm and GRACE [30]) and influence scores [17].
 Results
-Table 4 shows that Rank-Surprisal Ratio consistently exhibits strong correlation with post-training reasoning performance across all student models, achieving an average Spearman correlation of 0.86 and outperforming all alternative metrics.
+Table 4 shows that RSR consistently exhibits strong correlation with post-training reasoning performance across all student models, achieving an average Spearman correlation of 0.86 and outperforming all alternative metrics.
 These results indicate the effectiveness and practical value of RSR.
-In contrast, surprisal-based and rank-based metrics alone yield substantially weaker correlations (at most 0.59), highlighting the importance of capturing both informativeness and alignment through the rank-surprisal ratio.
+In contrast, surprisal-based and rank-based metrics alone yield substantially weaker correlations (at most 0.59).
+Analysis (e.g., § 3.2) suggests that they tend to emphasize high-likelihood trajectories while insufficiently capturing informativeness.
+This highlights the importance of jointly modeling informativeness and alignment via the rank-surprisal ratio.
 4.2 Ablation Study
@@
 The comparable correlations observed under reduced sample size and alternative hyperparameter settings (e.g., $r_{max}=500$) indicate that RSR is robust to both data scarcity and reasonable variations in $r_{max}$.
-Additional ablation results are in § C.3.
+Additional ablation results are in § E.2.
 5 Practical Applications
@@
 Model names are abbreviated as Q for Qwen and L for LLaMA.
-Additional results, including the GPQA evaluation and the No-Selection baseline, are provided in § C.4.
+Further results, including complete scores and additional GPQA evaluation, are provided in § F and § E.3.
 Experimental Setting
@@
 These results underscore the practical value of RSR for selecting effective trajectories prior to training.
-Additional trajectory selection results, including complete scores, further analysis, the No-Selection baseline, GPQA evaluation, and experiments under a reduced-teacher setting, are provided in § C.4.
+Complete scores are reported in § F.
+Additional trajectory selection results, including GPQA evaluation, No-Selection baseline, further analyses, and reduced-teacher experiments, are provided in § E.3.
 5.2 Teacher Selection
@@
 To better capture practical constraints, we consider a low-resource setting in which generating full training data for every candidate teacher model is either costly or infeasible.
-Instead, we sample a small set of 200 trajectories from each candidate teacher, evaluate them using different metrics, and select the teacher model based on the resulting dataset-level average score.
+Instead, we sample a small set of 200 trajectories from each candidate teacher, evaluate them using data-student suitability metrics, and select the teacher model based on the resulting dataset-level average score.
 We use 6 diverse teacher models (Deepseek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, and GPT-OSS-20B) as the candidate pool, avoiding consistently well-performing teachers to ensure a non-trivial selection task.
@@
 Prior work shows that stronger teachers do not necessarily yield better students, often due to capability gaps [50, 25] or off-policy data [5].
-Recent studies address this by better aligning teacher supervision with student behavior, achieving an implicit balance through approaches such as on-policy distillation [2], integration with reinforcement learning [26, 51], SFT optimization [44, 39], teaching assistants [28, 9], and interleaved sampling [40, 32].
+Recent studies address this by better aligning teacher supervision with student behavior, achieving an implicit balance through approaches such as on-policy distillation [2], self-distillation [49, 35], integration with reinforcement learning [26, 53], SFT optimization [45, 40], teaching assistants [28, 9], uncertainty-based filtering [51], and interleaved sampling [41, 32].
 By contrast, our work explicitly quantifies the trade-off between informativeness and alignment in distillation, enabling a principled identification of effective teacher data for a given student.
@@
 Recognizing that optimal reasoning data may vary across students, recent work explores student-specific trajectory selection strategies [30, 19, 17].
-Our work also studies student-specific data selection, but differs from prior work by measuring data–student suitability from the perspective of informative alignment and by conducting more comprehensive evaluations across a wider range of teacher models.
-7 Conclusion
-In this paper, we study data–student suitability in reasoning distillation and propose Rank-Surprisal Ratio (RSR), a simple metric for identifying suitable reasoning trajectories for a given student.
-Motivated by our analysis, RSR jointly captures a trajectory’s informativeness and alignment with the student’s behavior, favoring trajectories with low absolute probability but relatively high-ranked tokens.
-Experiments across diverse teacher–student pairs show that RSR strongly correlates with post-training performance and consistently outperforms existing metrics.
-We further demonstrate its effectiveness in both trajectory and teacher selection.
-Overall, our results highlight informative alignment as a promising direction for reasoning distillation.
+Our work also studies student-specific data selection, but differs from prior work by measuring data-student suitability from the perspective of informative alignment and by conducting more comprehensive evaluations across a wider range of teacher models.
 7 Broader Applicability of RSR
@@
 Although our work focuses solely on reasoning tasks, RSR is not specifically designed for reasoning trajectories (CoTs) and can be applied to general text data.
-Exploring its behavior and effectiveness beyond reasoning settings remains an interesting direction for future work.
+Exploring its behavior and effectiveness beyond reasoning settings remains an interesting direction for future work, and we welcome efforts from the community in this direction.
 Subset Selection
@@
 This scenario is particularly relevant when only a single trajectory is available per problem, yet improved data efficiency is still desired.
-Although subset selection introduces additional confounding factors due to variation across problems—making it harder to isolate trajectory–student suitability effects—RSR may still help identify relatively more effective samples across different reasoning problems.
+Although subset selection introduces additional confounding factors due to variation across problems—making it harder to isolate trajectory-student suitability effects—RSR may still help identify relatively more effective samples across different reasoning problems.
 Preliminary experiments on internal data provide suggestive evidence for this possibility.
+8 Conclusion
+In this paper, we study data-student suitability in reasoning distillation and propose Rank-Surprisal Ratio (RSR), a simple metric for identifying suitable reasoning trajectories for a given student.
+Motivated by our analysis, RSR jointly captures a trajectory’s informativeness and alignment with the student’s behavior, favoring trajectories with low absolute probability but relatively high-ranked tokens.
+Experiments across diverse teacher-student pairs show that RSR strongly correlates with post-training performance and consistently outperforms existing metrics.
+We further demonstrate its effectiveness in both trajectory and teacher selection.
+Overall, our results highlight informative alignment as a promising direction for reasoning distillation.
 Limitations
@@
 A promising direction for future work is to use our metric to guide the rewriting or synthesis of reasoning trajectories, rather than selecting from a fixed pool.
-In addition, the derived metric takes a simple and intuitive form with a clear interpretation.
-However, it remains unclear whether it arises from deeper mathematical principles.
-We have not yet identified a suitable theoretical framework to analyze this aspect, which we leave for future investigation.
-Finally, due to resource constraints, we primarily focus on mathematical reasoning tasks with extensive controlled studies, while also including additional evaluation on GPQA.
-Extending our evaluation to other domains,such as code or additional forms of reasoning, would be valuable.
-However, since our analysis relies on large-scale trajectory generation and approximately 200 SFT experiments, such extensions would require substantially greater computational resources and are therefore left for future work.
+In addition, given the simple and interpretable form of the derived RSR metric, it is natural to ask whether it can be grounded in deeper theoretical principles.
+However, we have not yet identified a suitable theoretical framework to characterize RSR, which we leave for future investigation.
+Finally, due to resource constraints, we focus primarily on mathematical reasoning tasks to conduct extensive controlled studies.
+Although we perform additional evaluation on GPQA to assess the generalizability of RSR, we acknowledge that the current experiments still have limitations.
+Specifically, we have not yet applied RSR to newly generated trajectories beyond the existing mathematical reasoning data, nor to qualitatively different domains such as code or commonsense reasoning.
+Systematic cross-domain studies would be valuable, but they require large-scale trajectory regeneration and retraining, entailing substantial computational cost.
+Therefore, we focus on the current setting (following prior work) and leave such systematic extensions for future work.
 Appendix A Details of Experiments
@@
 Second, high-quality reasoning problems are relatively scarce, making further scaling often unrealistic in practice.
-Third, our distillation study incurs quadratic computational costs across teacher–student pairs.
+Third, our distillation study incurs quadratic computational costs across teacher-student pairs.
 Considering these factors, we find 5,000 problems to offer a good trade-off between representativeness and computational feasibility.
@@
 For example, trajectories generated by GPT-OSS models tend to be concise, whereas those from DeepSeek-R1 are generally more detailed.
-Specific trajectory examples from different teachers are available in our released datasets on Hugging Face and in our case study (§ C.5).
+Specific trajectory examples from different teachers are available in our released datasets on Hugging Face and in our case study (§ D.2).
 For the student models, we select five open-source pretrained base models from the Qwen and LLaMA families: Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B.
@@
 Our evaluated benchmarks, AIME’25, AIME’24, AMC’23, and MATH500, are widely used and span varying difficulty levels, assessing mathematical reasoning and multi-step problem-solving across diverse domains.
-We additionally conduct evaluation on GPQA-Diamond beyond mathematics, as described in § C.4.2.
+We additionally conduct evaluation on GPQA-Diamond beyond mathematics, as described in § E.3.2.
 We adopt the Acc@4 metric as the final score, which averages results over four independent evaluations per problem.
@@
 We use the Spearman correlation coefficient [48] because our analysis focuses on monotonic consistency rather than strict linear relationships between metrics and post-training performance.
-We also report Pearson correlation [7] results in § C.2.
+We also report Pearson correlation [7] results in § E.1.
 We report the absolute values of Spearman correlation coefficients in Table 4.
@@
 We compare the correlation strength of the simple dataset-level average of trajectory-level RSR, i.e., $\frac{1}{|\mathbf{X}|}\sum_{j=1}^{|\mathbf{X}|}\mathrm{RSR}(\mathbf{x}_{j})$, with that of the dataset-level RSR computed via surprisal-weighted averaging (Eq.
-9) in the additional ablation study (§ C.3).
+9) in the additional ablation study (§ E.2).
 Appendix B Details of Compared Metrics
 B.1 LLM-judged Quality
-We use Qwen3-32B-Instruct in a Non-Thinking configuration as an automatic judge to evaluate reasoning trajectories.
+We use Qwen-3-32B-Instruct in a Non-Thinking configuration as an automatic judge to evaluate reasoning trajectories.
 Under a fixed evaluation prompt, the judge produces a structured assessment at two levels.
@@
 $\mathrm{Avg\_Surp_{local}}(\mathbf{x})=\frac{1}{T}\sum_{k=1}^{T}-\log p_{\theta}(t_{k}\mid\mathbf{c}_{k}^{\text{local}})$ (11)
-Appendix C More Results and Analysis
-C.1 Computational Cost for Rank-Surprisal Ratio
+Appendix C Computational Cost of RSR
 The computation of Rank-Surprisal Ratio requires only a single forward pass through the student model, during which we collect each token’s surprisal and clipped rank from the model logits.
@@
 This computation does not introduce additional complexity beyond that inherent in the dimensionality of the model logits.
-Moreover, since RSR only depends on clipped ranks, the rank computation can be efficiently implemented via top-$k$ selection rather than full-vocabulary comparisons, reducing practical runtime and GPU memory usage.
-In practice, computing RSR over the 5,000-trajectory dataset with a context length of 32,768 using a 7B model typically takes under one hour on a single H200 GPU with FlashAttention-2, which is significantly cheaper than SFT.
-C.2 Additional Results for Correlation Analysis
-Metric Correlation Measure Student Models
-Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
-Rank-Surprisal Ratio Spearman 0.855 0.845 0.918 0.818 0.845 0.856
-Pearson 0.654 0.880 0.805 0.819 0.811 0.794
-Table 10: Comparison of Spearman and Pearson correlation results on Rank-Surprisal Ratio (absolute values).
-Table 10 reports both Spearman and Pearson correlations for Rank-Surprisal Ratio.
-The results show that RSR also exhibits strong Pearson correlation with post-training performance.
-The rationale for primarily using Spearman correlation rather than Pearson correlation is discussed in § A.7.
-Table 23, 24, 25, 26, and 27 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
-C.3 Additional Ablation Study
-Variants Avg. Corr. $\Delta$
-Rank-Surprisal Ratio $0.856$
-   Fixed student $0.785$ 0.071-0.071
-   Simple dataset-level average $0.838$ 0.018-0.018
-   Use Rank1.05 $0.845$ 0.011-0.011
-   Use Rank0.95 $0.787$ 0.069-0.069
-   Use Surprisal1.05 $0.847$ 0.009-0.009
-   Use Surprisal0.95 $0.873$ $0.017$
-Avg-Rank (clipped) $0.552$ 0.304-0.304
-Rank Minus Surprisal $0.585$ 0.271-0.271
-Rank-Entropy Ratio $0.764$ 0.092-0.092
-Table 11: Additional ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation.
-We conduct some additional ablation studies. Results are shown in Table 11.
-The “Fixed student” variant computes RSR using a fixed model (Qwen-3-14B) instead of the target student, and the resulting drop in correlation highlights the importance of student-specific estimation.
-The "Simple dataset-level average" variant computes the dataset-level RSR by simply averaging trajectory-level scores, as noted in § A.8.
-The slight degradation in correlation indicates that a simple average of trajectory-level RSR remains a robust estimator of dataset-level suitability, while the surprisal-weighted averaging scheme improves the reliability of the aggregated metric at both trajectory-level and dataset-level.
-The “Avg-Rank (clipped)” variant applies rank clipping with $r_{max}=100$ to Avg-Rank; however, it still exhibits a notable performance gap compared with RSR.
-This suggests that rank information alone is insufficient to yield strong correlation, underscoring the importance of jointly leveraging rank and surprisal.
-The “Rank Minus Surprisal” metric is computed by subtracting the surprisal from the (clipped) rank.
-Although it also models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance.
-"Rank–Entropy Ratio" is defined as the ratio between the average (clipped) token rank and the average token entropy, where entropy is loosely related to surprisal.
-While this metric achieves relatively strong correlation, its performance remains below that of RSR.
-We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to a power of 1 for both rank and surprisal), for example "use Rank1.05".
-The results indicate that our metric is robust to these variations and may yield higher correlation with hyperparameter tuning.
-Nevertheless, we use the default power-1 setting to keep the formulation simple.
-C.4 Additional Results for Trajectory Selection
-C.4.1 RSR-based Trajectory Selection Variants and No-Selection Baseline
-Variant Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
-Math Avg. Math Avg. Math Avg. Math Avg. Math Avg.
-Rank-Surprisal Ratiomin (5k) 78.6 28.5 53.2 61.4 34.8 51.3
-   With Correctness Filtering 77.5 27.9 52.3 60.7 34.9 50.7
-   Fewer Candidates per Teacher 77.6 27.8 52.6 60.9 33.8 50.5
-No Selection (55k combined) 72.7 41.5 54.4 58.9 35.3 52.5
-Table 12: Performance of RSR-based trajectory selection variants and a No-Selection baseline trained on a 55k combined trajectory dataset.
-"Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
-Although trajectory selection enables data-efficient training and offers substantial practical value under resource-constrained settings, we are also interested in how it compares with training on all available data without any selection when resources are sufficient.
-To this end, we construct a “No Selection” baseline.
-Unlike RSR, which selects 5k trajectories from different teachers, this baseline directly merges 11 trajectory datasets—one per teacher—into a unified training set comprising 55k samples.
-We do not use the full 165k dataset (i.e., all three trajectory datasets per teacher) for two reasons.
-First, the computational cost would be prohibitively high.
-Second, training on multiple highly similar trajectory datasets from the same teacher may reduce data diversity and increase the risk of overfitting.
-We consider the 55k setting to be sufficiently representative and more consistent with practical use cases.
-From Table 12, we observe that the 5k trajectories selected by RSRachieve post-training performance that is comparable to—or even exceeds—the 55k-sample “No Selection” baseline on most student models.
-An exception is LLaMA-3.1-8B, which may require larger-scale data to develop reasoning ability due to its relatively limited prior exposure to reasoning tasks.
-In contrast, for stronger students (Qwen-3-14B and Qwen-3-4B), the selected 5k trajectories significantly outperform the unfiltered 55k dataset.
-This suggests that a small amount of high-quality data is sufficient to effectively activate the reasoning capabilities of student models.
-Overall, these findings demonstrate that trajectory selection not only improves data efficiency but can also deliver superior performance with substantially fewer resources, further highlighting the practical value of RSR.
-Table 12 also presents trajectory selection results for two selection method variants based on Rank-Surprisal Ratio, offering further insights.
-For "With Correctness Filtering", we consider a combined setting that uses both RSR and verified correctness for trajectory selection.
-Specifically, for problems with verifiably correct trajectories, we first discard incorrect ones and then select among the remaining correct trajectories based on RSR.
-The results show no significant improvement over selecting trajectories solely based on RSR, particularly for larger student models.
-This suggests that correctness is sometimes less critical than overall data suitability.
-For "Fewer Candidates per Teacher", we evaluate an 11-to-1 setting in which each candidate pool contains 11 teacher trajectories (one per teacher), instead of the original 33-to-1 setting (three per teacher).
-This setting focuses on selecting the best trajectory for each problem across different teachers, rather than across multiple generations from the same teacher.
-The results are comparable, with a slight performance gap relative to the 33-to-1 setting, indicating that RSR remains effective even when each teacher provides only a single trajectory.
-These findings further suggest that while RSR captures suitability differences across generations from the same teacher, such differences are less pronounced than those across different teacher models.
-C.4.2 Additional Evaluation on GPQA and Full Results
-Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
-GPQA GPQA GPQA GPQA GPQA
-Random 48.5 26.8 35.4 43.4 22.2 35.3
-LLM-judged Qualitymax 53.5 27.3 35.4 45.5 21.2 36.6
-Rank-Surprisal Ratiomin 55.1 31.3 38.9 45.5 31.3 40.4
-Table 13: Comparison of different trajectory selection methods on the GPQA-Diamond benchmark across student models.
-To more comprehensively evaluate the impact of different trajectory selection methods on post-trained models’ reasoning capabilities beyond mathematical problems, we conduct additional evaluation on the GPQA-Diamond benchmark [34].
-GPQA-Diamond consists of 198 challenging multiple-choice questions spanning biology, physics, and chemistry.
-Table 13 summarizes the evaluation results of different trajectory selection methods on GPQA-Diamond.
-Although the results are less stable than those on mathematical benchmarks due to the out-of-domain nature of this evaluation, datasets selected by RSR still achieve the best overall post-training performance.
-This suggests that RSR can identify suitable trajectories that consistently improve student models’ general reasoning capabilities, even when training solely on mathematical problems.
-The complete trajectory selection results underlying Table 6 are presented in Table 18, 19, 20, 21, and 22.
-C.4.3 Analysis of Datasets Selected by RSR
-Selected Datasets Data Composition over Teacher Models Metrics
-R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR Length
-RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57 9363
-RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69 9939
-RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67 9845
-RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56 9419
-RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73 10169
-Table 14: Data composition and metric statistics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) for different student models.
-Model names are abbreviated as Q for Qwen and L for LLaMA.
-Table 14 shows the data composition of datasets selected by RSR across 11 teacher models.
-The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students.
-QwQ-32B is generally preferred across student models, consistent with its stable performance.
-For a clearer contrast in data composition among student models, we refer readers to § C.4.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed.
-We release the RSR-selected datasets for five student models on Hugging Face.
-The average RSR values of the selected datasets are also reported in Table 14.
-These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § C.2 for complete metrics), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
-C.4.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
-To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers.
-Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B.
-This teacher set is formed by combining the teachers used in Table 2 and 7.
-Selection Methods Qwen-3-14B Qwen-2.5-7B
-Math Avg. Math Avg.
-Random 72.8 47.3
-LLM-judged Qualitymax 74.4 48.3
-Rank-Surprisal Ratiomin 76.8 50.0
-Table 15: Comparison of different trajectory selection methods under a reduced-teacher setting. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
-The results are shown in Table 15.
-Datasets selected by RSR still achieve superior post-training reasoning performance compared with the baselines, demonstrating the effectiveness of our metric when only a limited number of teachers are available.
-We also observe that the performance gap narrows when selecting from seven teachers compared with eleven teachers, which is expected and suggests that a larger candidate space enables trajectory selection to more effectively identify high-quality training data.
-Moreover, Table 16 presents the data composition of datasets selected by RSR from the reduced set of seven teacher models.
-The distributions differ markedly for Qwen-3-14B and Qwen-2.5-7B: the former student model tends to favor teachers such as DeepSeek-R1 and Qwen-3-30B, whereas the latter student model shows a stronger preference for smaller models such as Qwen-3-8B.
-These results further demonstrate the effectiveness of RSR in selecting teacher trajectories that are well suited to specific student models.
-Selected Datasets Data Composition over Teacher Models
-Deepseek-R1 Qwen-3-235B Nemotron Qwen-3-30B Magistral GPT-OSS-20B Qwen-3-8B
-RSRmin on Qwen-3-14B 27.40% 21.00% 1.42% 28.84% 4.92% 0.24% 16.18%
-RSRmin on Qwen-2.5-7B 14.52% 8.60% 6.94% 20.76% 10.64% 0.14% 38.40%
-Table 16: Data composition of trajectory datasets selected by RSR under a reduced-teacher setting (see the setting in § C.4.4 and results in Table 15).
-C.5 What Does RSR Prefer? A Token-Level Case Study
+Moreover, since RSR only depends on clipped ranks, the rank computation can be efficiently implemented via top-$k$ selection rather than full-vocabulary comparisons (see our code on GitHub), reducing practical runtime and GPU memory usage.
+In practice, computing RSR over the 5,000-trajectory dataset with a context length of 32,768 using a 7B model typically takes under one hour on a single H200 GPU with FlashAttention-2 enabled, which is significantly cheaper than SFT.
+We also conduct a direct wall-clock time comparison for computing RSR, G-Norm, Influence Score, and LLM-judged Quality on a fixed dataset of 5,000 DeepSeek-R1 trajectories using the LLaMA-3.1-8B student model.
+As shown in Table 10, RSR incurs significantly lower total GPU-hours than the other metrics.
+This likely stems from the fact that RSR requires only a single forward pass through the student model, whereas the other methods involve gradient computations or text generation.
+Metrics #GPU Hours Total GPU-hours
+RSR 1 1 1
+G-Norm 2 3 6
+Influence Score 8 4 32
+LLM-judged Quality 8 1.5 12
+Table 10: Computational cost comparison of different metrics on a fixed set of 5,000 DeepSeek-R1 trajectories using the LLaMA-3.1-8B student model, evaluated on H200 GPUs.
+Appendix D What Does RSR Prefer? Statistical Analysis and Case Study
+D.1 Statistical Analysis of RSR-Selected Trajectories
+We compare the statistics of the RSR-selected trajectories with the aggregate statistics (mean/max/min) across source teacher datasets.
+From Table 11, the RSR-selected trajectories exhibit moderate token length and verified accuracy.
+This suggests that RSR is not driven solely by length or accuracy, but instead follows its own evaluation criterion.
+Additional analysis of datasets selected by RSR can be found in § E.3.3.
+Datasets Avg. Token Length Verified Accuracy
+RSR Selected 9845 81.19
+Mean over Teachers 8912 81.58
+Max. 13145 85.9
+Min. 2552 77.6
+Table 11: Statistics of RSR-selected trajectories and source teacher trajectories in terms of token length and verified accuracy.
+Qwen-2.5-7B is used as the student model.
+RSR follows its own evaluation criterion, independent of length or accuracy.
+D.2 Token-Level Case Study
 To better understand how Rank-Surprisal Ratio operates in practice and what kinds of trajectories it prefers, we conduct a token-level case study comparing trajectories generated by different teacher models.
@@
 These observations are consistent with the distillation results (Table 1) and the subsequent analysis presented in the main body of the paper.
-We also observe that token-level RSR may become numerically unstable when extremely low surprisal yields inflated RSR values (e.g., 100+).
+We also observe that token-level RSR may become numerically unstable when extremely low surprisal yields inflated RSR values (e.g., >100).
 This motivates our surprisal-weighted averaging for trajectory-level RSR (§ 3.4), which stabilizes the metric and improves its effectiveness.
@@
 Each token is color-coded according to its RSR value, where darker red indicates lower values (stronger preference under RSR).
-GPT-OSS-20B exhibits uncommon phrasing and high rank values; Nemotron-Super exhibits conventional reasoning flow and low surprisal; QwQ-32B strikes a more favorable balance between informativeness and alignment, yielding lower token-level RSR values.
+GPT-OSS-20B exhibits uncommon phrasing and high rank values; Nemotron-Super exhibits conventional reasoning flow and low surprisal; QwQ-32B strikes a more favorable balance between informativeness and alignment, yielding lower token-level RSR values and thus being more likely to be selected.
 Trajectory-level RSR is defined as the surprisal-weighted average of token-level RSR values with rank clipping (§ 3.4).
-Appendix D Others
-D.1 License for Artifacts and Data Consent
-All artifacts used in this paper are publicly available for academic research purposes, including AIME, AMC, MATH500, and NuminaMath.
-D.2 Data Statement
+Appendix E More Results and Analysis
+E.1 Additional Results for Correlation Analysis
+Table 13 reports both Spearman and Pearson correlations for Rank-Surprisal Ratio.
+The results show that RSR also exhibits strong Pearson correlation with post-training performance.
+The rationale for primarily using Spearman correlation rather than Pearson correlation is discussed in § A.7.
+Metric Correlation Measure Student Models
+Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
+Rank-Surprisal Ratio Spearman 0.855 0.845 0.918 0.818 0.845 0.856
+Pearson 0.654 0.880 0.805 0.819 0.811 0.794
+Table 13: Comparison of Spearman and Pearson correlation results on Rank-Surprisal Ratio (absolute values).
+E.2 Additional Ablation Study
+We conduct additional ablation studies with more design variants of RSR. Results are shown in Table 14.
+Variants Avg. Corr. $\Delta$
+Rank-Surprisal Ratio $0.856$
+   Fixed student $0.785$ 0.071-0.071
+   Simple dataset-level average $0.838$ 0.018-0.018
+   Use Rank1.05 $0.845$ 0.011-0.011
+   Use Rank0.95 $0.787$ 0.069-0.069
+   Use Surprisal1.05 $0.847$ 0.009-0.009
+   Use Surprisal0.95 $0.873$ $0.017$
+Avg-Rank (clipped) $0.552$ 0.304-0.304
+Avg-Entropy $0.495$ 0.361-0.361
+Surprisal-Weighted Sum of Rank $0.604$ 0.252-0.252
+Rank-Weighted Sum of Surprisal $0.584$ 0.272-0.272
+Rank Minus Surprisal $0.585$ 0.271-0.271
+Rank Times Surprisal $0.538$ 0.318-0.318
+Rank-Entropy Ratio $0.764$ 0.092-0.092
+Table 14: Additional ablation study for Rank-Surprisal Ratio. $\Delta$ denotes the change in average correlation. All ranks are clipped at $r_{max}=100$.
+The “Fixed student” variant computes RSR using a fixed model (Qwen-3-14B) instead of the target student, and the resulting drop in correlation highlights the importance of student-specific estimation.
+The "Simple dataset-level average" variant computes the dataset-level RSR by simply averaging trajectory-level scores, as noted in § A.8.
+The slight degradation in correlation indicates that a simple average of trajectory-level RSR remains a robust estimator of dataset-level suitability, while the surprisal-weighted averaging scheme improves the reliability of the aggregated metric at both trajectory-level and dataset-level.
+The “Avg-Rank (clipped)” variant applies rank clipping with $r_{max}=100$ to Avg-Rank (see Table 4); however, it still exhibits a notable performance gap compared with RSR.
+This suggests that rank information alone is insufficient to yield strong correlation, underscoring the importance of jointly leveraging rank and surprisal.
+Similarly, other metrics, such as average entropy and weighted-sum combinations, yield substantially lower correlation than RSR.
+Our analysis suggests that these alternatives still tend to emphasize high-likelihood trajectories while insufficiently capturing informativeness.
+The “Rank Minus Surprisal” metric is computed by subtracting surprisal from the (clipped) rank.
+Although it models the relationship between rank and surprisal, this simple subtraction does not align well with post-training performance; similarly, “Rank Times Surprisal” also shows poor alignment.
+The “Rank-Entropy Ratio” is defined as the ratio between the average (clipped) token rank and the average token entropy, where entropy is loosely related to surprisal.
+Notably, it exhibits a smaller performance drop, suggesting that ratio-based formulations better balance alignment and informativeness.
+We also experiment with different exponent choices when computing Rank-Surprisal Ratio (with the default setting corresponding to a power of 1 for both rank and surprisal), for example "use Rank1.05".
+The results indicate that our metric is robust to these variations and may yield higher correlation with hyperparameter tuning.
+Nevertheless, we use the default power-1 setting to keep the formulation simple.
+E.3 Additional Results for Trajectory Selection
+E.3.1 RSR-based Trajectory Selection Variants and No-Selection Baseline
+Although trajectory selection enables data-efficient training and offers substantial practical value under resource-constrained settings, we are also interested in how it compares with training on all available data without any selection when resources are sufficient.
+To this end, we construct a “No Selection” baseline.
+Unlike RSR, which selects 5k trajectories from different teachers, this baseline directly merges 11 trajectory datasets—one per teacher—into a unified training set comprising 55k samples.
+We do not use the full 165k dataset (i.e., all three trajectory datasets per teacher) for two reasons.
+First, the computational cost would be prohibitively high.
+Second, training on multiple highly similar trajectory datasets from the same teacher may reduce data diversity and increase the risk of overfitting.
+We consider the 55k setting to be sufficiently representative and more consistent with practical use cases.
+Variant Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
+Math Avg. Math Avg. Math Avg. Math Avg. Math Avg.
+Rank-Surprisal Ratiomin (5k) 78.6 28.5 53.2 61.4 34.8 51.3
+   With Correctness Filtering 77.5 27.9 52.3 60.7 34.9 50.7
+   Fewer Candidates per Teacher 77.6 27.8 52.6 60.9 33.8 50.5
+No Selection (55k combined) 72.7 41.5 54.2 58.9 35.3 52.5
+Table 15: Performance of RSR-based trajectory selection variants and a No-Selection baseline trained on a 55k combined trajectory dataset.
+"Math Avg." denotes the average performance over AIME’24, AIME’25, AMC’23, and MATH500.
+From Table 15, we observe that the 5k trajectories selected by RSR achieve post-training performance that is comparable to—or even exceeds—the 55k-sample “No Selection” baseline on most student models.
+An exception is LLaMA-3.1-8B, which may require larger-scale data to develop reasoning ability due to its relatively limited prior exposure to reasoning tasks.
+In contrast, for stronger students (Qwen-3-14B and Qwen-3-4B), the selected 5k trajectories significantly outperform the unfiltered 55k dataset.
+This suggests that a small amount of high-quality data is sufficient to effectively activate the reasoning capabilities of student models.
+Overall, these findings demonstrate that trajectory selection not only improves data efficiency but can also deliver superior performance with substantially fewer resources, further highlighting the practical value of RSR.
+Table 15 also presents trajectory selection results for two selection method variants based on RSR, offering further insights.
+For "With Correctness Filtering", we consider a combined setting that uses both RSR and verified correctness for trajectory selection.
+Specifically, for problems with verifiably correct trajectories, we first discard incorrect ones and then select among the remaining correct trajectories based on RSR.
+The results show no significant improvement over selecting trajectories solely based on RSR, particularly for larger student models.
+This suggests that correctness is sometimes less critical than overall data suitability.
+For "Fewer Candidates per Teacher", we evaluate an 11-to-1 setting in which each candidate pool contains 11 teacher trajectories (one per teacher), instead of the original 33-to-1 setting (three per teacher).
+This setting focuses on selecting the best trajectory for each problem across different teachers, rather than across multiple generations from the same teacher.
+The results are comparable, with a slight performance gap relative to the 33-to-1 setting, indicating that RSR remains effective even when each teacher provides only a single trajectory.
+The results also suggest that while RSR captures suitability differences across generations from the same teacher, such differences are less pronounced than those across different teacher models.
+E.3.2 Additional Evaluation on GPQA
+To more comprehensively evaluate the impact of different trajectory selection methods on post-trained models’ reasoning capabilities beyond mathematical problems, we conduct additional evaluation on the GPQA-Diamond benchmark [34].
+GPQA-Diamond consists of 198 challenging multiple-choice questions spanning biology, physics, and chemistry.
+Selection Methods Qwen-3-14B LLaMA-3.1-8B Qwen-2.5-7B Qwen-3-4B Qwen-2.5-3B Average
+GPQA GPQA GPQA GPQA GPQA
+Random 48.5 26.8 35.4 43.4 22.2 35.3
+LLM-judged Qualitymax 53.5 27.3 35.4 45.5 21.2 36.6
+Rank-Surprisal Ratiomin 55.1 31.3 38.9 45.5 31.3 40.4
+Table 16: Comparison of different trajectory selection methods on the GPQA-Diamond benchmark across student models. We report Acc@4 as the evaluation metric.
+Table 16 summarizes the evaluation results of different trajectory selection methods on GPQA-Diamond.
+Although the results are less stable than those on mathematical benchmarks due to the out-of-domain nature of this evaluation, datasets selected by RSR still achieve the best overall post-training performance.
+This suggests that RSR can identify suitable trajectories that consistently improve student models’ general reasoning capabilities, even when training solely on mathematical problems.
+Complete tables combining mathematical and GPQA evaluation results can be found in § F.
+E.3.3 Analysis of Datasets Selected by RSR
+Table 17 shows the data composition of datasets selected by RSR across 11 teacher models.
+The resulting distributions vary across student models, demonstrating the metric’s ability to select different teacher trajectories tailored to different students.
+QwQ-32B is generally preferred across student models, consistent with its stable performance.
+For a clearer contrast in data composition among student models, we refer readers to § E.3.4, where trajectory selection is performed with fewer teachers and consistently strong teachers such as QwQ-32B are removed.
+We release the RSR-selected datasets for five student models on Hugging Face.
+The average RSR values of the selected datasets are also reported in Table 17.
+These datasets consistently achieve substantially lower RSR values than the teacher trajectory datasets (see § F), validating that our selection procedure effectively identifies trajectories with low RSR for each problem.
+Selected Datasets Data Composition over Teacher Models Metrics
+R1 Q3-235B GPT-120B Nemotron QwQ Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B RSR Length
+RSRmin on Q3-14B 6.2% 4.5% 0.1% 0.1% 67.3% 6.7% 1.4% 0.0% 2.9% 2.1% 8.6% 2.57 9363
+RSRmin on L3.1-8B 5.5% 1.0% 0.0% 3.7% 48.8% 3.9% 9.1% 0.1% 0.2% 21.3% 6.5% 2.69 9939
+RSRmin on Q2.5-7B 4.1% 1.0% 0.7% 1.7% 55.7% 5.3% 5.9% 0.2% 0.8% 17.6% 6.9% 2.67 9845
+RSRmin on Q3-4B 2.7% 2.2% 0.0% 0.3% 76.6% 4.1% 3.1% 0.0% 0.6% 4.5% 6.1% 2.56 9419
+RSRmin on Q2.5-3B 2.4% 0.8% 0.0% 2.4% 45.1% 4.0% 12.2% 0.0% 0.2% 26.0% 6.6% 2.73 10169
+Table 17: Data composition and metric statistics of trajectory datasets selected by RSR in the trajectory selection experiments (§ 5.1) for different student models.
+Model names are abbreviated as Q for Qwen and L for LLaMA.
+E.3.4 Additional Trajectory Selection Experiments with Fewer Teacher Models
+To better reflect practical scenarios in which only a few teachers’ trajectories are available and generally suitable teachers may be absent, we conduct additional experiments that select trajectories from a reduced set of teachers.
+Specifically, we select trajectories from candidates generated by seven teachers: DeepSeek-R1, Qwen-3-235B-Thinking, Nemotron-Super, Qwen-3-30B-Thinking, Magistral-Small, GPT-OSS-20B, and Qwen-3-8B.
+This teacher set is formed by combining the representative teachers used in Table 2 and § 5.2.
+The results are shown in Table 18.
+Datasets selected by RSR still achieve superior post-training reasoning performance compared with the baselines, demonstrating the effectiveness of our metric when only a limited number of teachers are available.
+We also observe that the performance gap narrows when selecting from seven teachers compared with eleven teachers, which is expected and suggests that a larger candidate space enables trajectory selection to more effectively identify high-quality training data.
+Selection Methods Qwen-3-14B Qwen-2.5-7B
+Math Avg. Math Avg.
+Random 72.8 47.3
+LLM-judged Qualitymax 74.4 48.3
+Rank-Surprisal Ratiomin 76.8 50.0
+Table 18: Comparison of different trajectory selection methods under a reduced-teacher setting. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
+Moreover, Table 19 presents the data composition of datasets selected by RSR from the reduced set of seven teacher models.
+The distributions differ markedly for Qwen-3-14B and Qwen-2.5-7B: the former student model tends to favor teachers such as DeepSeek-R1 and Qwen-3-30B, whereas the latter student model shows a stronger preference for smaller models such as Qwen-3-8B.
+These results further demonstrate the effectiveness of RSR in selecting teacher trajectories that are well suited to specific student models.
+Selected Datasets Data Composition over Teacher Models
+Deepseek-R1 Qwen-3-235B Nemotron Qwen-3-30B Magistral GPT-OSS-20B Qwen-3-8B
+RSRmin on Qwen-3-14B 27.40% 21.00% 1.42% 28.84% 4.92% 0.24% 16.18%
+RSRmin on Qwen-2.5-7B 14.52% 8.60% 6.94% 20.76% 10.64% 0.14% 38.40%
+Table 19: Data composition of trajectory datasets selected by RSR under a reduced-teacher setting (see the setting in § E.3.4 and results in Table 18).
+Appendix F Complete Results Tables
+Table 25, 26, 27, 28, and 29 present the full metric assessment results across different teacher trajectory datasets on Qwen-3-14B, LLaMA-3.1-8B, Qwen-2.5-7B, Qwen-3-4B, and Qwen-2.5-3B, respectively.
+The complete trajectory selection evaluation results underlying Table 6 are presented in Table 20, 21, 22, 23, and 24.
+Appendix G Others
+G.1 License for Artifacts and Data Consent
+All artifacts used in this paper are publicly available for academic research purposes, including AIME, AMC, MATH500, GPQA-Diamond and NuminaMath.
+G.2 Data Statement
 The training datasets consist solely of mathematics problems and solutions and contain no offensive content or personal information.
-D.3 AI Assistant Usage Statement
+G.3 AI Assistant Usage Statement
 We used ChatGPT for writing refinement and minor coding assistance. AI assistants were not involved in research innovation, and all core contributions were developed solely by the authors.
@@
 Rank-Surprisal Ratiomin 67.5 59.2 93.1 94.6 78.6 55.1
-Table 18: Full post-training evaluation results for trajectory selection on Qwen-3-14B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
+Table 20: Full post-training reasoning evaluation results for trajectory selection on Qwen-3-14B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
 Selection Methods LLaMA-3.1-8B
@@
 Rank-Surprisal Ratiomin 5.0 8.3 36.9 63.6 28.5 31.3
-Table 19: Full post-training evaluation results for trajectory selection on LLaMA-3.1-8B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
+Table 21: Full post-training reasoning evaluation results for trajectory selection on LLaMA-3.1-8B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
 Selection Methods Qwen-2.5-7B
@@
 Rank-Surprisal Ratiomin 29.2 25.8 71.3 86.6 53.2 38.9
-Table 20: Full post-training evaluation results for trajectory selection on Qwen-2.5-7B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
+Table 22: Full post-training reasoning evaluation results for trajectory selection on Qwen-2.5-7B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
 Selection Methods Qwen-3-4B
@@
 Rank-Surprisal Ratiomin 44.2 35.0 77.5 88.8 61.4 45.5
-Table 21: Full post-training evaluation results for trajectory selection on Qwen-3-4B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
+Table 23: Full post-training reasoning evaluation results for trajectory selection on Qwen-3-4B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
 Selection Methods Qwen-2.5-3B
@@
 Rank-Surprisal Ratiomin 11.7 11.7 45.0 70.8 34.8 31.3
-Table 22: Full post-training evaluation results for trajectory selection on Qwen-2.5-3B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
+Table 24: Full post-training reasoning evaluation results for trajectory selection on Qwen-2.5-3B. "Math Avg." denotes the average over AIME’24, AIME’25, AMC’23, and MATH500.
 Metrics Deepseek-R1 Q3-235B GPT-120B Nemotron QwQ-32B Q3-30B Magistral GPT-20B Phi-4 Q3-8B Q3-4B
@@
 RSR 2.925 2.940 3.527 3.352 2.673 2.923 3.302 3.645 3.360 3.003 2.918
+Post-Training Performance 77.1 71.8 66.7 72.2 77.4 77.2 68.8 69.5 54.1 74.6 76.8
 Table 25: Full metric assessment results on Qwen-3-14B. Model names are abbreviated with Q for Qwen.
@@
 RSR 2.996 3.044 3.971 3.016 2.818 2.965 3.020 4.038 3.633 2.882 2.945
+Post-Training Performance 28.1 22.0 15.2 23.7 27.1 26.7 22.8 17.9 14.5 26.5 28.2
 Table 26: Full metric assessment results on LLaMA-3.1-8B. Model names are abbreviated with Q for Qwen.
@@
 RSR 3.002 3.023 3.686 3.086 2.779 2.951 3.091 3.827 3.468 2.888 2.940
+Post-Training Performance 47.3 45.0 40.7 48.3 52.0 50.0 47.6 42.7 35.2 52.0 51.8
 Table 27: Full metric assessment results on Qwen-2.5-7B. Model names are abbreviated with Q for Qwen.
@@
 RSR 2.958 2.955 3.794 3.216 2.649 2.917 3.160 3.888 3.588 2.919 2.928
+Post-Training Performance 55.8 53.4 47.9 56.4 61.2 58.8 52.2 48.4 40.2 61.2 61.9
 Table 28: Full metric assessment results on Qwen-3-4B. Model names are abbreviated with Q for Qwen.
@@
 RSR 3.095 3.103 3.944 3.107 2.860 3.012 3.050 4.037 3.622 2.911 2.994
+Post-Training Performance 29.6 26.4 22.9 33.0 33.0 31.2 30.6 24.4 18.2 34.2 33.3
 Table 29: Full metric assessment results on Qwen-2.5-3B. Model names are abbreviated with Q for Qwen.
 Evaluation prompt for LLM-judged quality assessment
-Table 28: Evaluation prompt for LLM-judged quality assessment.
+Table 30: Evaluation prompt for LLM-judged quality assessment as a baseline metric.